### v1

# Chapter 0 Strong CP problem, theta term and QCD topological properties

Abstract
[Abstract] In this chapter we introduce the $\theta$-dependence and the topological properties of QCD, features of the strongly interacting sector which give rise to the strong CP problem in the more general context of the Standard Model of particle physics. We discuss the analytical approaches that can be used to obtain qualitative, or in some cases quantitative, information on the $\theta$-dependence of QCD and QCD-like models, discussing their range of validity and comparing their predictions with the numerical results obtained by means of lattice simulations.
keywords:
[…]
Objectives
  • In Sect. 2 we introduce $\theta$-dependence in a didactic but not completely standard way, avoiding the use of the semiclassical approximation; we then show how $\theta$-dependence leads to the strong CP problem.
  • In Sect. 3 we describe some general features of $\theta$-dependence, discussing in particular the predictions that can be obtained by using the semiclassical approximation, the expansion in the number of colors, and chiral perturbation theory. These predictions are then compared with the numerical results obtained using the lattice formulation.
1 Introduction
Quantum Chromodynamics (QCD) is the non-abelian gauge theory describing the fundamental strong interactions among quarks and gluons within the Standard Model of particle physics. While high-energy processes characterized by a large momentum transfer can be reliably studied using perturbation theory, this is not possible in the low energy regime, where the theory is strongly coupled. This regime is dominated by non-perturbative effects, that are responsible for several intriguing theoretical features which have fundamental implications for particle physics phenomenology. The topics of this chapter—QCD topology, $\theta$-dependence and the strong CP problem—fall under this category.
The QCD $\theta$-dependence can be seen as a consequence of the non-trivial topological features of the space of gluon fields, since configurations can be classified based on the value of their topological charge:
$Q=\frac{g^{2}}{16\pi^{2}}\epsilon_{\mu\nu\rho\sigma}\int\mathrm{d}^{4}x\,\mathrm{Tr}\left[F^{\mu\nu}(x)F^{\rho\sigma}(x)\right]\,.$ (1)
As the name suggests, this quantity has an inherently topological meaning: it can only assume integer values for gauge configurations with finite Euclidean action, and it is impossible to change its value via a continuous field deformation. $\theta$-dependence is usually introduced by using semiclassical arguments, which lead to interpret $Q$ as the number of windings of the gluon field around the gauge group at infinite distance, but a different approach is also possible, which will be discussed in the next section. Being the topological charge $Q$ dimensionless and gauge-invariant, there is no theoretical obstruction in coupling it to the ordinary QCD action via a new dimensionless parameter $\theta$:
$\mathcal{S}_{{\scriptscriptstyle{\mathrm{QCD}}}}\longrightarrow\mathcal{S}_{{\scriptscriptstyle{\mathrm{QCD}}}}+\theta Q\ .$ (2)
While this topological $\theta$ term would be completely irrelevant at the classical level (being it the integral of a four-divergence), it has actually physical consequences in the quantum theory. Indeed the path-integral which is expected to non-perturbatively define QCD as a Quantum Field Theory can be decomposed into a sum over all the topological sectors (i. e. the subsets of the gauge field space with fixed values of the topological charge) and the introduction of a $\theta$ term changes the relative weight of the different sectors, thus introducing a non-trivial dependence on this parameter in all physical observables.
One of the most fundamental consequences of $\theta$-dependence for Standard Model phenomenology is the explicit violation of the charge conjugation and parity (CP) symmetry it introduces in the strong sector, being $Q$ CP-odd by virtue of the presence of the Levi-Civita tensor. So far, however, no strong CP violation has ever been observed in experiments, leading to a very stringent upper bound on $|\theta|$, which is set to be of the order of $\sim 10^{-9}-10^{-10}$ by the non-observation of the neutron electric dipole moment. This introduces a fine-tuning problem on $\theta$ known as strong CP problem, one of the most relevant open problems of the Standard Model.
One of the most promising proposed solution to the strong CP problem is the Peccei–Quinn axion, an hypothetical particle whose peculiar coupling to gluons dynamically suppresses strong CP violations. Axion detection is currently the target of several major experimental collaborations, and has motivated plenty of theoretical studies about axion dynamics. In this respect, accurate predictions about QCD $\theta$-dependence turn out to be crucial to describe the axion effective potential. While in some regimes it is possible to obtain reliable predictions about $\theta$-dependence using suitable approximations, one has in general to resort to numerical Monte Carlo simulations of the lattice-discretized theory to systematically address this task from first-principles and in a fully non-perturbative way.
The goal of this chapter is to pedagogically introduce the reader to $\theta$-dependence in QCD and to the formulation of the strong-CP problem (Section 2), providing a compact compilation of state-of-the-art non-perturbative results about QCD topology and $\theta$-dependence obtained by means of both analytical tools and numerical methods (Section 3).
2 $\theta$-dependence and the strong CP problem
1 Why $\theta$-dependence
The $\theta$-dependence of four dimensional gauge theories is typically introduced using the semiclassical approximation and instantons, see, e. g., ColemanBook ; ShifmanBook ; Weinberg2Book ; ZinnJustinBook . A canonical approach can however also be used, which emphasizes the role of locality and of the topology of the gauge group (as opposed to the topology of field configurations with finite Euclidean action), clarifying that $\theta$-dependence is a physical property of the theory independent of the semiclassical approximation. Such an approach has been developed in Jackiw:1979ur ; Jackiw85 , and its influence can be seen, e. g., in the presentations of WeinbergEBook and, particularly, of SmilgaBook .
Before addressing the case of four-dimensional gauge theories, let us start by discussing in some detail what is arguably the technically simplest system characterized by $\theta$-dependence, i. e., a quantum particle moving on a circumference. This system, despite its extreme simplicity, displays several key features which have analogues in four-dimensional gauge theories Jackiw:1979ur ; SmilgaBook ; SchulmanBook ; Bietenholz:1997kr ; Gaiotto:2017yup ; Bonati:2017woi . If we denote by $\phi$ the angle on the circumference, the Hamiltonian of a quantum particle of mass $m$ moving on a circumference of radius $R$ is
$H=\frac{1}{2mR^{2}}p_{\phi}^{2}+V(\phi)=-\frac{\hbar^{2}}{2mR^{2}}\frac{\partial^{2}}{\partial\phi^{2}}+V(\phi)\ ,$ (3)
where $p_{\phi}=-\mathrm{i}\hbar\frac{\partial}{\partial\phi}$ and $V(\phi)$ is an external field satisfying the geometrical constraint $V(2\pi)=V(0)$. We moreover assume the external potential to be parity invariant, i.e. $V(\phi)=V(2\pi-\phi)$, and real. The Hamiltonian $H$, to represent a well defined self-adjoint operator, needs to be supplemented by boundary conditions on the interval $0,2\pi)$. The periodic boundary conditions $\psi(0)=\psi(2\pi)$ are typically imposed to the wavefunction $\psi(\phi)$, but more general conditions can also be adopted.
A physically sensible way to find more general boundary conditions is to note that, if we extend the angle $\phi$ from $0,2\pi)$ to the whole $\mathbb{R}$, the system is characterized by a $2\pi$ translation invariance in $\phi$. If we denote by $U$ the $2\pi$ translation operator $U\psi(\phi)=\psi(\phi+2\pi)$, this operator is thus an unitary operator which implements a symmetry of the system, hence it commutes with the Hamiltonian $H$, and $U$ and $H$ can be diagonalized simultaneously. Considering the eigenvalue $\mathrm{e}^{\mathrm{i}\theta}$ (the notation is not accidental, as we will show in a moment) of the unitary operator $U$, the corresponding eigenvector $|\psi_{\theta}\rangle$ satisfies $U|\psi_{\theta}\rangle=\mathrm{e}^{\mathrm{i}\theta}|\psi_{\theta}\rangle$, which corresponds to periodic up-to-a-phase boundary conditions for the wave function
$\psi_{\theta}(2\pi)=\mathrm{e}^{\mathrm{i}\theta}\psi_{\theta}(0)\ .$ (4)
Since $U$ commutes with the Hamiltonian, eigenspaces corresponding to different eigenvalues of $U$ will not interfere with each other in the dynamics of the system, and we can thus use Eq. (4) as boundary condition for the differential operator in Eq. (3). This is just the starting point needed to prove the Bloch theorem in solid state physics, see, e. g., AshcroftMerminBook , and Eq. (4) is in fact the most general boundary condition which makes the differential operator $p_{\phi}=-\mathrm{i}\hbar\frac{\partial}{\partial\phi}$ self-adjoint on the $0,2\pi)$ interval, see, e. g., ReedSimon2Book , or Bonneau:1999zq for an elementary discussion. Already at this level we can see that, since for values of $\theta$ different from 0 and $\pi$ we have $\mathrm{e}^{\mathrm{i}\theta}\neq\mathrm{e}^{-\mathrm{i}\theta}$, the boundary condition in Eq. (4) explicitly breaks the symmetries under parity ($\mathrm{P}\psi(\phi)=\psi(2\pi-\phi)$) and time reversal ($\mathrm{T}\psi(\phi)=\psi^{*}(\phi)$) that characterize the classical system.
Path integration provides a nice way of moving $\theta$-dependence from the boundary condition to the action, which will be useful also in the case of four dimensional gauge theories. The propagator of the theory obtained without fixing the value of $\theta$ can be explicitly written as
$\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle=\int_{0}^{2\pi}\mathrm{d}\theta\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta)}(t_{f}-t_{i})/\hbar}\ ,$ (5)
where $\psi_{\theta,n}$ is the $n$-th eigenfunction of the problem with boundary condition Eq. (4), and using path integration (see, e. g., SchulmanBook ; FeynmanHibbsBook ; ZinnJustinPIBook ) this expression can be rewritten in the form (we always consider $\phi\in\mathbb{R}$)
$\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle=\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)\ ,\quad L=p_{\phi}\dot{\phi}-H=\frac{1}{2}mR^{2}\dot{\phi}^{2}-V(\phi)\ .$ (6)
Since $\mathrm{e}^{\mathrm{i}\theta}$ is the eigenvalue of the operator $U$ which has been diagonalized simultaneously with $H$ we also have
$_{\theta^{\prime}}\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle_{\theta}=2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta)}_{n}(t_{f}-t_{i})/\hbar}=2\pi\delta(\theta-\theta^{\prime})\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}\ .$ (7)
We can now use the identity
$\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}=2\pi\delta(\theta-\theta^{\prime})$ (8)
to select the value of $\theta$ we are interested in: using Eq. (4) and Eq. (6) we have
$\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=$ (9)
$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f}+2\pi Q)\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\langle\phi_{f}+2\pi Q,t_{f}|\phi_{i},t_{i}\rangle=$
$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{\phi(t_{i})=\phi_{i}}^{\phi(t_{f})=\phi_{f}+2\pi Q}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)=\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}\left(L(\phi,\dot{\phi})-\hbar\frac{\theta}{2\pi}\dot{\phi}\right)\mathrm{d}t\right)\ ,$
and we thus obtain the Lagrangian
$L_{\theta}=\frac{1}{2}mR^{2}\dot{\phi}^{2}-V(\phi)-\hbar\frac{\theta}{2\pi}\dot{\phi}\ .$ (10)
The last term of the Lagrangian $L_{\theta}$ has the typical form of a $\theta$ term: it is a term of quantum origin (as signaled by the explicit presence of $\hbar$ in the Lagrangian) which is a total derivative with respect to time, hence it has no effect on the classical equation of motion of the system, still it has non trivial effects in the quantum theory. To verify the last statement, i. e. the effect of the $\theta$ term in the quantum theory, there are (at least) two ways. The first, more direct, one consists in explicitly computing the energy spectrum of the model in the free case $V=0$: it is indeed trivial to verify that in this case the energy spectrum explicitly depends on $\theta$, and it is given by
$E_{n}^{(\theta)}=\frac{\hbar^{2}}{2mR^{2}}\left(n+\frac{\theta}{2\pi}\right)^{2}\ ,\quad n\in\mathbb{Z}\ .$ (11)
The second method uses the path-integral formulation: the $\theta$ term contributes to the propagator from the state $|\phi_{i}\rangle$ at time $t_{i}$ to the state $|\phi_{f}\rangle$ at time $t_{f}$ with
$\exp\left(-\frac{\mathrm{i}\theta}{2\pi}\int_{t_{i}}^{\,t_{f}}\dot{\phi}\mathrm{d}t\right)=\exp\left(-\frac{\mathrm{i}\theta}{2\pi}[\phi_{f}(t_{f})-\phi_{i}(t_{i})]\right)\ .$ (12)
The effect of the $\theta$ term is thus of topological origin, as can be seen by noting that, for fixed $\phi_{i}(t_{i})$ and $\phi_{f}(t_{f})$, the contributions of two different paths have a relative weight $\exp(-\mathrm{i}\theta\Delta Q)$, where $\Delta Q$ is the difference between the winding numbers along the circumference of the two paths. For the particular case of the quantum particle moving on a circumference we can also note that $\theta$ enters the Lagrangian $L_{\theta}$ in the same way as a constant vector potential directed along the $\phi$ direction:
$L_{A_{\phi}}=\frac{1}{2}mR^{2}\dot{\phi}^{2}-V(\phi)-\frac{e}{c}R\dot{\phi}A_{\phi}\ ,$ (13)
and by introducing the magnetic flux $\Phi_{B}=2\pi RA_{\phi}$ we can write
$\frac{\theta}{2\pi}=\frac{\Phi_{B}}{\Phi_{0}}\ ,$ (14)
where we defined the elementary magnetic flux $\Phi_{0}=\frac{2\pi\hbar c}{e}$. For this simple model $\theta$-dependence is thus just a manifestation of the Ahronov-Bohm effect, see, e. g., BallentineBook ; KonishiPaffutiBook .
We are now ready to discuss the case of four dimensional gauge theories, considering for the sake of the simplicity the case of the gauge group SU($N$) and using natural units. The Lagrangian density of the pure glue theory is
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}^{a}F^{a\,\mu\nu}$ (15)
where $F^{a}_{\mu\nu}=\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{a}+gf^{abc}A_{\mu}^{b}A_{\nu}^{c}$ is the $a$-th component of field strength tensor, $a=1,\ldots,N^{2}-1$ denotes the color component, $f^{abc}$ are the SU($N$) structure constants, and $g$ is the gauge coupling. The possible presence of matter fields does not significantly affect the following discussion, except for some details that will be signaled in due time. The first step required to start the canonical quantization programme is to write the Hamiltonian density corresponding to the Lagrangian density in Eq. (15) (something which is in fact required also in the path integral construction), and we immediately face a well known problem: the variable $\dot{A}^{0}$ does not enter Eq. (15), so we can not define its conjugate momentum $P_{0}$. The theory of constrained Hamiltonian systems has been developed, starting from Dirac DiracBook , to deal with singular Lagrangians, i. e. those cases in which the Legendre transformation between momenta and velocities is not well defined. A simple account of the basic ideas can be found in SmilgaBook , while more details are presented, e. g., in Weinberg1Book ; HansonReggeTeitelboimBook ; HenneauxTeitelboimBook , and the outcome of this analysis is that it is legitimate to consider the spatial component $A_{j}^{a}$, $j=1,2,3$ and their conjugate momenta $P_{j}^{a}$ as dynamical variables subject to the constraint given by the equation of motion obtained by varying $\mathcal{L}$ with respect to $A^{0}$. In particular it is simple to verify that $P_{j}^{a}=F_{0j}^{a}$ and that the constraint can be written in the form
$G^{a}\equiv\partial_{j}P_{j}^{a}+gf^{abc}A_{j}^{b}P_{j}^{c}=0\ .$ (16)
Note that, since the structure constants are antisymmetric in all indices, there is no problem related to the operator ordering in the second term of the previous expression.
To understand the physical meaning of the constraint in Eq. (16) we have to remember how gauge transformations acts on the gauge field: if we introduce for the sake of the simplicity the matrix notation $A_{\mu}=A_{\mu}^{a}T^{a}$ (where $T^{a}$ are the Hermiatian generators of the fundamental SU($N$) representation, normalized by $\mathrm{Tr}(T^{a}T^{b})=\frac{1}{2}\delta^{ab}$) a finite gauge transformation $\Omega(x)$ acts on the gauge field as follows
$A_{\mu}\to A_{\mu}^{\Omega}=\Omega(x)A_{\mu}\Omega^{{\dagger}}(x)-\frac{\mathrm{i}}{g}[\partial_{\mu}\Omega(x)]\Omega^{{\dagger}}(x)\ ,$ (17)
and in particular, by writing $\Omega(x)=\mathrm{e}^{\mathrm{i}gT^{a}\chi^{a}(x)}\approx 1+\mathrm{i}gT^{a}\chi^{a}(x)$, the change of the spatial components of the gauge field under an infinitesimal gauge transformation is
$\delta A_{j}^{a}=\partial_{i}\chi^{a}-gf^{abc}\chi^{b}A_{j}^{c}\ .$ (18)
All gauge transformations that will be used in the following will be assumed to be local, i. e. $\Omega(\mathbf{x})\to 1$ sufficiently fast as $|\mathbf{x}|\to\infty$ (we denote by $\mathbf{x}$ the spatial components of $x$, at fixed time). Infinitesimal gauge transformations, in particular, have to satisfy $\chi^{a}(\mathbf{x})\to 0$ as $|\textbf{x}|\to\infty$. If we now denote by $\Psi[\mathbf{A}]$ the wavefunction of the system (with A denoting collectively the spatial components $A_{j}(\mathbf{x})$ for a fixed time), the variation of $\Psi[\mathbf{A}]$ due to an infinitesimal gauge transformation is (integrating by part and using $P^{a\,j}(x)=-\mathrm{i}\frac{\delta}{\delta A_{j}^{a}(x)}$)
$\Psi[\mathbf{A}+\delta\mathbf{A}]-\Psi[\mathbf{A}]\simeq\int\frac{\delta\Psi[\mathbf{A}]}{\delta A_{j}^{a}}\delta A_{j}^{a}\mathrm{d}^{3}\mathbf{x}=-\int\left(\chi^{a}\partial_{j}\frac{\delta}{\delta A_{j}^{a}}+gf^{abc}\chi^{a}A_{j}^{b}\frac{\delta}{\delta A_{j}^{c}}\right)\Psi[\mathbf{A}]\mathrm{d}^{3}\mathbf{x}\propto\int\chi^{a}(\mathbf{x})G^{a}\Psi[\mathbf{A}]\mathrm{d}^{3}\mathbf{x}\ .$ (19)
Since $\chi^{a}(\mathbf{x})$ approaches zero for large x, the previous integral is well defined, and we see that the constraint in Eq. (16) implements gauge invariance under infinitesimal local gauge transformations. If matter is coupled to the gauge field, a further term appears in Eq. (16), related to the variation with respect to $A^{0}$ of the matter part of the Lagrangian density. The resulting constraint once again implements gauge invariance, this time under infinitesimal local gauge transformations of both gauge and matter fields.
Not all finite local gauge transformations can however be continuously connected to the identity by infinitesimal transformations. Since $\Omega(\mathbf{x})$ approaches the identity as $|\textbf{x}|\to\infty$, $\Omega(\mathbf{x})$ effectively defines a map from the 3-sphere $S^{3}$ (which is the one-point-compactification of $\mathbb{R}^{3}$, obtained by adding the point at infinity) to the gauge group, and for SU($N$) the third homotopy group is nontrivial: $\pi_{3}(\mathrm{SU}(N))=\mathbb{Z}$. This can be made more concrete by introducing the quantity
$N(\Omega)=\frac{1}{24\pi^{2}}{\epsilon}^{ijk}\int\mathrm{Tr}\left[\Omega^{-1}(\partial_{i}\Omega)\Omega^{-1}(\partial_{j}\Omega)\Omega^{-1}(\partial_{k}\Omega)\right]\mathrm{d}^{3}\mathbf{x}\ .$ (20)
It is indeed possible to show (see, e. g., Weinberg2Book ; WeinbergEBook ) that $N(\Omega)$ is the winding number of the mapping $\Omega(\mathbf{x})$ from $S^{3}$ to SU($N$). In particular $N(\Omega)$ is invariant under infinitesimal changes of $\Omega$, and $N(\Omega_{1}\Omega_{2})=N(\Omega_{1})+N(\Omega_{2})$. It is also possible to show, by explicit computation, that the winding number of the transformation
$\Omega_{1}=\exp\left(\mathrm{i}\frac{x_{j}\sigma_{j}}{|\mathbf{x}|}f(|\textbf{x}|)\right)\ ,\$(21)
where $f$ is a monotonic function with $f(0)=-\pi$ and $f(\infty)=0$, is $N(\Omega_{1})=1$, hence we have an explicit example of a gauge transformation with nonvanishing winding number. Such topologically nontrivial gauge transformations are often called “large” gauge transformations, as opposed to the “small” ones, which are continuously connected to the identity.
The large gauge transformation $\Omega_{1}$ defined above is the equivalent of the $2\pi$ rotation for the quantum particle on a circumference. If we denote by $U$ the unitary operator implementing the gauge transformation $\Omega_{1}$ in the Hilbert space of the theory, since the theory is gauge invariant $U$ commutes with the Hamiltonian, but the constraint in Eq. (16) is not sufficient to fix $U=1$. The best we can do is to simultaneously diagonalize $U$ and the Hamiltonian, once again fixing the eigenvalue $\mathrm{e}^{\mathrm{i}\theta}$ of the unitary operator $U$. In terms of the wavefunction $\Psi(\mathbf{A})$ this amounts to impose $\Psi(\mathbf{A}^{\Omega_{1}})=\mathrm{e}^{\mathrm{i}\theta}\Psi(\mathbf{A})$, which implies, given the properties of $N(\Omega)$ and the invariance under infinitesimal transformations of $\Psi$, the more general
$\Psi(\mathbf{A}^{\Omega})=\mathrm{e}^{\mathrm{i}\theta N(\Omega)}\Psi(\mathbf{A})\ .$ (22)
We are now almost ready to adapt the reasoning carried out in Eq. (9) to the case of four dimensional gauge theories. The last ingredient needed is a relation between the winding number $N(\Omega)$ and four dimensional integrals. To this purpose we introduce the (non gauge invariant) Chern-Simons current
$j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\epsilon^{\,\mu\alpha\beta\gamma}\mathrm{Tr}\left(F_{\alpha\beta}A_{\gamma}+\frac{2\mathrm{i}g}{3}A_{\alpha}A_{\beta}A_{\gamma}\right)\ ,$ (23)
its gauge invariant four-divergence (usually called topological charge density)
$q\equiv\partial_{\mu}j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\mathrm{Tr}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)=\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (24)
and its integral on a fixed time surface
$W(\mathbf{A})=\int j_{CS}^{0}(t,\textbf{x})\mathrm{d}^{3}\textbf{x}\ .$ (25)
Using the four dimensional divergence theorem and assuming the spatial components of the Chern-Simons current to be trivial at spatial infinity (e. g. due to some fixed boundary conditions) we have
$\int q(x)\,\mathrm{d}^{4}x=\int(\partial_{0}j^{0}_{CS}+\partial_{k}j^{k}_{CS})\mathrm{d}^{4}x=\int_{t_{i}}^{t_{f}}\mathrm{d}t\,\partial_{0}W(\mathbf{A})=W(\textbf{A}_{f})-W(\textbf{A}_{i})\ ,$ (26)
moreover under the gauge transformation $\mathbf{A}\to\textbf{A}^{\Omega}$ it is possible to check by direct computation Jackiw85 that $W(\mathbf{A})$ transforms as follows
$W(\mathbf{A}^{\Omega})=W(\textbf{A})-N(\Omega)\ .$ (27)
We can now repeat almost verbatim Eq. (9):
$\left.\langle\mathbf{A}_{f},t_{f}|\textbf{A}_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu(\theta-\theta^{\prime})}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=$ (28)
$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\langle\mathbf{A}_{f}^{\Omega_{1}^{\nu}},t_{f}|\textbf{A}_{i},t_{i}\rangle=$
$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}^{\Omega_{1}^{\nu}}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\mathcal{L}\right)\propto\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\left[\mathcal{L}+\theta q\right]\right)\ ,$
where in the last step we used the gauge invariance of $\mathcal{L}$ under time dependent gauge transformations and
$\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}\mathbf{x}\,q(t,\mathbf{x})=W(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})-W(\mathbf{A}_{i})=W(\mathbf{A}_{f})-W(\mathbf{A}_{i})-\nu\ .$ (29)
2 The strong CP problem
The Lagrangian density of QCD, taking into account also the matter fields and the $\theta$ term, is thus
$\mathcal{L}=-\frac{1}{4}F^{a}_{\mu\nu}F^{a\,\mu\nu}+\sum_{{\scriptscriptstyle{\mathrm{f}}}}\bar{\psi}_{{\scriptscriptstyle{\mathrm{f}}}}(\mathrm{i}\gamma_{\mu}D^{\mu}-m_{{\scriptscriptstyle{\mathrm{f}}}})\psi_{{\scriptscriptstyle{\mathrm{f}}}}+\theta\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (30)
where $D_{\mu}=\partial_{\mu}-igA_{\mu}^{a}T^{a}$ is the covariant derivative in the fundamental representation of the gauge group and “f” is a flavor index. It is well known that the first two terms of this Lagrangian density (just like in Quantum Electrodynamics) are invariant under both parity (P) and charge conjugation (C), however the $\theta$ term explicitly breaks P and CP for generic values of the $\theta$ angle, and this is essentially due to the presence of the Levi-Civita totally antisymmetric tensor. If we introduce the three dimensional chromoelectric and chromomagnetic fields in analogy with classical electromagnetism we have indeed that $q\propto\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, and since the chromomagnetic field is even under parity transformation we see that the $\theta$ term is odd under both P and CP. Invariance under parity and CP is clearly recovered for $\theta=0$, but also for $\theta=\pi$: since $\theta$ is defined up to $2\pi$, we have indeed $\pi\sim-\pi$. Note that at $\theta=\pi$ the discrete symmetries are however expected to be spontaneously broken, since at $\theta=\pi$ a level crossing happens, just like in Eq. (11) (see also the following discussion on $\theta$-dependence in chiral perturbation theory in Sec. 2 and Witten:1980sp ; Gaiotto:2017yup ).
Parity violations in strong interactions are (if they exist at all) very small, so it is meaningful to perform an expansion in $\theta$, which parametrizes the only term in the Lagrangian density which explicitly breaks P and CP symmetries. Violations of these symmetries in physical observables are thus expected to be linear in $\theta$ with good accuracy. The neutron electric dipole moment $d_{\rm n}$ has been identified as one of the most sensitive observables that can be used to identify violation of P and CP in strong interactions, and using dimensional analysis one gets the estimate
$|d_{\rm n}|\sim|\theta|\,e\,\frac{m_{\pi}^{2}}{m^{3}_{\rm n}}\ ,$ (31)
[…]
$d_{\rm n}=(-2.9\pm 0.9)\cdot 10^{-3}\,\theta\,e\,\mathrm{fm}\ .$ (32)
[…]
$|d_{\rm n}|=(0.0\pm 1.1_{\rm stat}\pm 0.2_{\rm syst})\cdot 10^{-13}\,e\,\mathrm{fm}$ (33)
gives the upper bound $|\theta|\lesssim 10^{-10}$ for the value of the $\theta$ parameter. So why not fixing directly $\theta=0$ in $\mathcal{L}_{\theta}$? If we only considered QCD, independently of the other sectors of the Standard Model, this would indeed be possible: it would be enough to impose from the beginning local gauge invariance, renormalizability, and invariance under parity to obtain $\mathcal{L}_{\theta}$ with $\theta=0$.
To discuss why this is not possible when considering the whole Standard Model we have to introduce one last ingredient: the chiral anomaly Adler:1969gk ; Bell:1969ts . If we consider a four dimensional gauge theory coupled to a single flavor of quark, the chiral rotation
$\psi\to\mathrm{e}^{\mathrm{i}\alpha\gamma^{5}}\psi\ ,\quad\bar{\psi}\to\bar{\psi}\mathrm{e}^{\mathrm{i}\alpha\gamma^{5}}$ (34)
can be shown to be equivalent to a complex rotation of the mass and to a shift of the $\theta$ angle (see, e. g., Weinberg2Book ):
$m\bar{\psi}\psi\to\bar{\psi}m^{\prime}\frac{1+\gamma^{5}}{2}\psi+\bar{\psi}m^{\prime*}\frac{1-\gamma^{5}}{2}\psi\ ,\quad m^{\prime}=m\mathrm{e}^{2\mathrm{i}\alpha}\ ,\quad\theta\to\theta+2\alpha\ .$ (35)
This relation implies that we can reabsorb $\theta$ in the mass term at the expense of using the complex mass $m\mathrm{e}^{-i\theta}$. Note that this does not change the invariance properties of the theory, since a complex mass is not invariant under P and CP transformation, but gives us the possibility of studying $\theta$-dependence using chiral perturbation theory if light quarks are present (and this is the rationale behind Eq. (31), since $m_{\pi}^{2}\propto m$). An immediate consequence of this fact is that there is no $\theta$-dependence at all if $m=0$, since in this case we can shift $\theta$ at will without affecting the mass term of the Lagrangian. When both the $\theta$ term and a complex mass are present, we can define the physical $\theta$ angle as the combination $\theta_{phys}=\theta-\arg m$, which is invariant under chiral transformations and corresponds to the value of the $\theta$ angle when the fermion mass is real and positive. The previous discussion can be generalized to the case of a generic number of flavors with mass term
$\bar{\psi}M\frac{1+\gamma^{5}}{2}\psi+\bar{\psi}M^{{\dagger}}\frac{1-\gamma^{5}}{2}\psi\ ,$ (36)
where $M$ is a $N_{f}\times N_{f}$ matrix and the physical $\theta$ angle is in this case given by
$\theta_{phys}=\theta-\arg\det M\ .$ (37)
The previous statement regarding the upper limit of the $\theta$ angle should thus more precisely be written as $|\theta_{phys}|\lesssim 10^{-10}$, and the presence of $\theta_{phys}$ shows why we can not simply fix to zero the $\theta$ angle in the Standard Model: the mass matrix originates from the electroweak sector, and since in the electroweak sector P and CP are not conserved, there is no way to fix $\theta_{phys}=0$ using symmetry arguments. The fact that $|\theta_{phys}|$ is so small or, equivalently, the absence of P and CP violations in the strong sector of the Standard Model, constitutes the so called strong CP problem. As noted before there is no $\theta$-dependence at all if at least one quark is massless ($\arg\det M$ is not defined if $M$ is singular). In the past, a massless up quarks has thus been suggested as a simple solution of the strong CP problem, however this is today known to be inconsistent with experimental and lattice QCD results FlavourLatticeAveragingGroupFLAG:2021npn ; Sanfilippo:2015era ; ParticleDataGroup:2024cfk (see also Alexandrou:2020bkd ).
3 General aspects of $\theta$-dependence
As we argued above, within the general framework of the Standard Model one cannot exclude a priori a $\theta$ term in the QCD Lagrangian. At the same time, the dependence of various properties of strong interactions on $\theta$ plays a significant role by itself, which is relevant to many phenomenological and theoretical issues, like the interpretation of some particular features of the hadronic spectrum, or the characterization of the various phases which are found when considering Yang–Mills theories or QCD at finite temperature. Moreover, one of the proposed solutions to the strong CP problem is the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur , which invokes the presence of a new particle, the axion, which has also become a popular candidate for dark matter: since the axion couples to QCD like a $\theta$ term, the study of $\theta$-dependence in QCD turns out to be fundamental in order to predict some properties of this dark matter candidate.
The purpose of this Section is to go more in detail into these interesting aspects of $\theta$-dependence, and to review the knowledge that we have achieved so far. It should be clear that, being linked to the existence of large field fluctuations with non-trivial topology, the problem is non-perturbative, hence the present knowledge is based either on general non-perturbative theoretical approaches to Yang–Mills theories, like, e. g., the large-$N$ expansion, or on numerical computations based on the lattice formulation.
The discussion about $\theta$-dependence is best formulated within the Euclidean path-integral formulation of the theory, which provides the framework for the lattice QCD numerical approach, and is also the natural setting for the finite temperature theory. Therefore, in order to fix ideas, we consider the path-integral formulation of the QCD thermal partition function which, up to an overall multiplicative normalization constant, reads:
[…]
where $\mathcal{S}_{{\scriptscriptstyle{\mathrm{E}}}}=\int\mathrm{d}^{4}x\,\mathcal{L}_{{\scriptscriptstyle{\mathrm{E}}}}$ is the Euclidean QCD action, $Q=\int\mathrm{d}^{4}x\,q(x)$, and
$\mathcal{L}_{{\scriptscriptstyle{\mathrm{E}}}}=\frac{1}{4}F^{a}_{\mu\nu}F^{a}_{\mu\nu}+\sum_{{\scriptscriptstyle{\mathrm{f}}}}\bar{\psi}_{{\scriptscriptstyle{\mathrm{f}}}}(\gamma^{{\scriptscriptstyle{\mathrm{E}}}}_{\mu}D_{\mu}+m_{{\scriptscriptstyle{\mathrm{f}}}})\psi_{{\scriptscriptstyle{\mathrm{f}}}}\ ;\ \ \ \ \ \ \ q(x)=\frac{g^{2}}{16\pi^{2}}F^{a}_{\mu\nu}\tilde{F}^{a}_{\mu\nu}\ ;\ \ \ \ \ \ \ \tilde{F}^{a}_{\mu\nu}=\frac{1}{2}\epsilon_{\,\mu\nu\alpha\beta}F^{a}_{\alpha\beta}\,$, (39)
where we have introduced the dual field strength tensor $\tilde{F}^{a}_{\mu\nu}$, while $\mathcal{S}_{{\scriptscriptstyle{\mathrm{E}}}}^{{\scriptscriptstyle{(\mathrm{YM})}}}$ refers to the pure gauge part of $\mathcal{S}_{{\scriptscriptstyle{\mathrm{E}}}}$. The Euclidean Lagrangian is defined in terms of the Wick rotated fields, $t\to-\mathrm{i}\tau$ where $\tau$ is the Euclidean time, and of the Euclidean gamma matrices ($\gamma_{0}^{{\scriptscriptstyle{\mathrm{E}}}}=\gamma^{0}\ ,\ \gamma_{j}^{{\scriptscriptstyle{\mathrm{E}}}}=-\mathrm{i}\gamma^{j}$). Space-time integration is assumed to be compactified in the Euclidean time direction, with a compactification extension $\beta_{T}=1/T$, where $T$ is the physical temperature, and periodic/antiperiodic boundary conditions in the time direction for bosonic/fermionic fields, in order to reproduce thermal conditions.
[…]
It is clear from Eq. (38) that, in the Euclidean formulation, the $\theta$ term gives a purely imaginary contribution to the action, making it complex. This is usually known as then sign problem, since it ruins the positivity of the Euclidean path-integral measure, thus undermining the Monte Carlo approach at its very basis. A similar problem appears in other cases, for instance, the fermion determinants themselves becomes complex after adding chemical potentials coupled to quark numbers. In this case, its origin can be traced back to the fact that the $\theta$ term involves an odd number of time derivatives (contrary to the standard pure gauge term), hence the appearance of the extra factor $\mathrm{i}$”.
That also underlies one of the exact results about $\theta$-dependence Vafa:1983tf ; Vafa:1984xg , the so-called Vafa–Witten theorem. Let us consider the $\theta$-dependent part of the free energy density (vacuum energy density at $T=0$):
$F(T,\theta)=-\frac{1}{\mathcal{V}}\log\left[\frac{Z(T,\theta)}{Z(T,0)}\right]$ (40)
where $\mathcal{V}$ is the 4-dimensional volume. The extra factor $\mathrm{i}$” of the Euclidean $\theta$ term implies that the ratio of partion functions can be viewed as the expectation value of a pure phase factor, $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$, which is then bounded by 1 from above, meaning that $F(T,\theta)\geq 0$ with a minimum in $\theta=0$ [^1] The same property holds for the quantum particle moving on a circumference discussed above: in that case, since a non-zero $\theta$ can be interpreted as the presence of a magnetic flux, the property implies that the system is diamagnetic. That rules out the possibility of spontaneous P breaking at $\theta=0$ Vafa:1984xg , hence $F(T,\theta)$ is an even function of $\theta$ with period $2\pi$ (because $Q\in\mathbb{Z}$) and analytic around $\theta=0$. Therefore, $F(T,\theta)$ can be written as a Taylor series around $\theta=0$, containing only even powers. A common parametrization is the following:
$F(T,\theta)=\frac{1}{2}\chi\theta^{2}\left(1+\sum_{n\,=\,1}^{\infty}b_{2n}\theta^{2n}\right)\,.$ (41)
Eq. (40) and the relation $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$ permit to rewrite the expansion coefficients in terms of expectation values computed at $\theta=0$
$\chi=\frac{1}{\mathcal{V}}\braket{Q^{2}}\ ;\ \ \ \ \ b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\frac{\braket{Q^{2n}}_{c}}{\braket{Q^{2}}}$ (42)
where $\braket{\cdot}_{c}$ stands for a connected expectation value, e. g., $\braket{Q^{4}}_{c}=\braket{Q^{4}}-3\braket{Q^{2}}^{2}$.
[…]
As we have discussed above, the problem why $\theta$ is so small does not find a solution within the Standard Model, given that the possibility of a massless up quark is ruled out. An attractive solution is provided by the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma , which invokes an extension of the Standard Model with an extra axial U(1) global symmetry, the U(1)PQ Peccei–Quinn symmetry. Such symmetry is supposed to be spontaneously broken, implying the possible emergence of a massless Nambu-Goldstone boson, the axion, but also anomalously broken at the same time, leading to a coupling of the axion field $a(x)$ to the topological charge density operator $q(x)$.
One can built various theories implementing such features, see e. g., Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq for early proposals and DiLuzio:2020wdo for a recent review, however, at the level of a low energy effective theory, they all correspond to the addition to the QCD Lagrangian of a massless pseudoscalar with derivative couplings plus a modification of the $\theta$ term as follows:
$\theta\,q(x)\ \ \to\ \ \left(\frac{a}{f_{a}}+\theta\right)q(x)$ (43)
where $f_{a}$ is the axion decay constant, which is supposed to be very large compared to the Standard Model scales. The key point of the mechanism is that, in this way, we have an effective $\theta_{{\scriptscriptstyle{\mathrm{eff}}}}=\theta+a/f_{a}$, which now participates to dynamics. In particular, the shift invariance $a(x)\to a(x)+\alpha$ that would be present if U(1)PQ were exact is explicitly broken by the anomalous coupling to $q(x)$, so that the axion field, hence $\theta_{{\scriptscriptstyle{\mathrm{eff}}}}$, will select the value which minimizes the vacuum energy density. By virtue of the Vafa–Witten theorem discussed above, such value is indeed $\theta_{{\scriptscriptstyle{\mathrm{eff}}}}=0$, which thus solves the CP-problem.
The axion serves multiple scopes: apart from solving the CP-problem, it is also an interesting dark matter candidate Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . The interesting point is that, because of its coupling to QCD, $\theta$-dependence fixes much of its physical properties. For instance, the second order expansion of the vacuum energy density around the minimum at $\theta_{{\scriptscriptstyle{\mathrm{eff}}}}=0$ can be interpreted as an effective axion mass term, leading to a relation between the axion mass $m_{a}$ and the topological susceptibility in QCD, $m^{2}_{a}=\chi/f_{a}^{2}$, while the higher order coefficients $b_{2n}$ can be related to the axion self couplings.
In the following, we will first review the available predictions and results on $\theta$-dependence regarding the low temperature phase of SU($N$) gauge theories, with or without fermions, which is characterized by relevant phenomena such as color confinement and chiral symmetry breaking, then exposing predictions and results concerning the high temperature phase, showing how $\theta$-dependence changes drastically at the transition between the two.
Before doing that, we will briefly review one of the simplest model which can be used to predict the form of the $\theta$-dependence, known as the Dilute Instanton Gas Approximation (DIGA) model. Non-abelian gauge theories are characterized by the existence of classical solutions to the equations of motion, which are either selfdual or anti-selfdual, i. e., $\tilde{F}_{\mu\nu}=\pm F_{\mu\nu}$, with finite action and non-zero topological charge $Q$. The most elementary of such solutions, with $Q=\pm 1$, are known as instantons and anti-instantons and carry an action $8\pi/g^{2}$, where $g$ is the gauge coupling: their field strength tensor is non-zero only over a limited portion of space and time (hence the name), each solution being characterized by a typical extension radius $\rho$. The reader will find an extensive treatment about instantons elsewhere in this Volume, as well as in various classical papers, reviews and textbooks (see, e. g., Refs. Callan:1977gz ; Gross:1980br ; Schafer:1996wv ; ColemanBook ).
The DIGA model represents a typical semiclassical approach, which assumes that the dominant contribution to the functional integral can be found by integrating fluctuations around single (anti)instantons. As we will discuss later on, such approach is expected to be reliable only at high $T$, while it fails at low $T$. There is one good reason to discuss it at this stage: it clarifies a key point, i. e., that for each phase of Yang–Mills theories, the functional dependence of $F(T,\theta)$ is tightly related to the distinctive topological degrees of freedom which are relevant in that phase.
1 A semiclassical model: the Dilute Instanton Gas Approximation
As anticipated above, the DIGA model builds around the hypothesis that the gauge field configurations giving the most relevant contribution to the functional integral $Z(T,\theta)$ are those around (i. e., differing by small fluctuations) the classical configurations obtained by superposing certain numbers of instantons (I) and anti-instanton (A) solutions. The numbers $n_{{\scriptscriptstyle{\mathrm{I}}}}/n_{{\scriptscriptstyle{\mathrm{A}}}}$ of I / A solutions for each configuration is arbitrary, however they are assumed to be dilute enough, i. e., with typical distances from each other much greater than their sizes, so that the total Euclidean action can be written as a sum of independent terms, and the functional integration of fluctuations around each I / A is independent of all the others. Under such assumptions, and using the fact that the topological charge of each configuration is $Q=n_{{\scriptscriptstyle{\mathrm{I}}}}-n_{{\scriptscriptstyle{\mathrm{A}}}}$, the thermal partition function can be written as
$Z(T,\theta)=\sum_{n_{{\scriptscriptstyle{\mathrm{I}}}}\,=\,0}^{\infty}\sum_{n_{{\scriptscriptstyle{\mathrm{A}}}}\,=\,0}^{\infty}\frac{1}{n_{{\scriptscriptstyle{\mathrm{I}}}}!n_{{\scriptscriptstyle{\mathrm{A}}}}!}Z_{1}^{n_{{\scriptscriptstyle{\mathrm{I}}}}}Z_{1}^{n_{{\scriptscriptstyle{\mathrm{A}}}}}\mathrm{e}^{\mathrm{i}\theta(n_{{\scriptscriptstyle{\mathrm{I}}}}-n_{{\scriptscriptstyle{\mathrm{A}}}})}=\left(\sum_{n_{{\scriptscriptstyle{\mathrm{I}}}}\,=\,0}^{\infty}\frac{Z_{1}^{n_{{\scriptscriptstyle{\mathrm{I}}}}}}{n_{{\scriptscriptstyle{\mathrm{I}}}}!}\mathrm{e}^{\mathrm{i}\theta n_{{\scriptscriptstyle{\mathrm{I}}}}}\right)\left(\sum_{n_{{\scriptscriptstyle{\mathrm{A}}}}\,=\,0}^{\infty}\frac{Z_{1}^{n_{{\scriptscriptstyle{\mathrm{A}}}}}}{n_{{\scriptscriptstyle{\mathrm{A}}}}!}\mathrm{e}^{-\mathrm{i}\theta n_{{\scriptscriptstyle{\mathrm{A}}}}}\right)=\exp\left\{\mathrm{e}^{\mathrm{i}\theta}Z_{1}\right\}\exp\left\{\mathrm{e}^{-\mathrm{i}\theta}Z_{1}\right\}=\exp\left\{2\cos(\theta)\,Z_{1}\right\}.$ (44)
where $Z_{1}$ is the single instanton functional integral, obtained by integrating fluctuations around the single I or A solution, which is actually a function of the temperature, $Z_{1}(T)$. The previous computation is in fact identical to the one needed to write the classical grand-canonical partition function of two different species of identical and non-interacting particles, I and A, with $Z_{1}$ the one-particle partition function; at $\theta=0$ one has a Poissonian distribution for both $n_{{\scriptscriptstyle{\mathrm{I}}}}$ and $n_{{\scriptscriptstyle{\mathrm{A}}}}$, which is a manifestation of diluteness, like for a perfect gas.
Using the definition of the free energy in Eq. (40), as well as its Taylor expansion representation in Eq. (41), one easily finds:
$F(T,\theta)=-\frac{1}{\mathcal{V}}\log\left[\frac{Z(T,\theta)}{Z(T,0)}\right]=\frac{2Z_{1}(T)}{\mathcal{V}}\left[1-\cos(\theta)\right]=\chi(T)\left[1-\cos(\theta)\right]\,.$ (45)
It is interesting to notice that the sole hypothesis of diluteness fixes $F(T,\theta)$, up to an overall pre-factor corresponding to the topological susceptibility $\chi(T)={2Z_{1}(T)}/{\mathcal{V}}$. In particular, one has a well defined and $T$-independent prediction for the $b_{2n}$ coefficients:
$b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\,$, (46)
which means, for instance, $b_{2}=-1/12$.
Let us now outline the main facts regarding the computation of $Z_{1}$, which can be carried out using perturbation theory on an instanton background (see, e. g., Refs. Gross:1980br ; Pisarski:1980md ; Boccaletti:2020mxu for more details). The result contains a contribution proportional to the exponential of the single instanton action, $\exp(-8\pi^{2}/g^{2}(\rho))$, where $g(\rho)$ is the running coupling computed at the scale of the instanton size $\rho$, so that the contribution increases/decreases, by asymptotic freedom, as $\rho$ increases/decreases; such contribution, however, must be integrated over the instanton size distribution which, for $N$ colors and $N_{{\scriptscriptstyle{\mathrm{f}}}}$ flavours turns out to be proportional to $\rho^{\frac{11}{3}N-\frac{2}{3}N_{{\scriptscriptstyle{\mathrm{f}}}}-5}$, which for $N=3$ diverges for large $\rho$ if $N_{{\scriptscriptstyle{\mathrm{f}}}}<9$. This fact signals the failure of the DIGA model: while it assumes diluteness, it is dominated by instantons of arbitrarily large size, which completely spoil diluteness.
However, the outcome changes if large size instantons are suppressed by other means: this is exactly what happens at finite temperature $T$, which acts as an infrared cutoff by limiting $\rho\lesssim 1/T$, i. e., by the extension of the compactified Euclidean temporal direction. Therefore one expects that, at least for large enough temperatures (how large, is not known a priori) the model should work. Moreover, for asymptotically large temperatures, one has the following predictions for the topological susceptibility (assuming $N_{{\scriptscriptstyle{\mathrm{f}}}}$ light flavours of equal mass $m$):
$\chi(T)\propto m^{N_{{\scriptscriptstyle{\mathrm{f}}}}}T^{4-\frac{11}{3}N-\frac{1}{3}N_{{\scriptscriptstyle{\mathrm{f}}}}}\,.$ (47)
Let us stress that the functional dependence for $F(T,\theta)$ predicted by Eq. (45), which is just based on diluteness, could start to be valid at temperatures lower than those at which the asymptotic behaviour of Eq. (47), which is the result of a 1-loop computation, sets in.
2 Low temperature: large-$N$ and chiral scaling
The topological properties of QCD in the low temperature phase are strictly entangled with hadron phenomenology, in particular with features of the hadron spectrum linked to the chiral anomaly. The most famous example is the role played by topology in explaining the mass of the $\eta^{\prime}$ meson Witten:1978bc ; Witten:1979vv ; Veneziano:1979ec ; Kawarabayashi:1980dp ; Witten:1980sp ; DiVecchia:1980yfw . The U(1)A symmetry that is present in chiral limit of the classical QCD Lagrangian can not be a quantum symmetry realized à la Wigner, since no parity doublet is found in the hadron spectrum. Moreover, as first pointed out by S. Weinberg Weinberg:1975ui , the mass of the singlet pseudoscalar meson $\eta^{\prime}$ should be smaller than $\sqrt{3}\,m_{\pi}\simeq 240$ MeV if the U(1)A symmetry were spontaneously broken, a bound which is largely violated by the actual mass, $m_{\eta^{\prime}}\simeq 958$ MeV. This is the so-called U(1)A puzzle. While it became soon clear that the bound violation must be related to the explicit breaking of U(1)A by the chiral anomaly tHooft:1976rip , the exact mechanism leading to the phenomenological value of $m_{\eta^{\prime}}$ is not obvious.
A useful approach to clarify the origin of the $\eta^{\prime}$ mass is to consider the limit of QCD known as the ’t Hooft large-$N$ limit tHooft:1973alw , in which we assume $N\to\infty$ at fixed $N_{\scriptscriptstyle{\mathrm{f}}}$, keeping the ’t Hooft coupling, $\lambda=g^{2}N$, fixed; the physical interpretation of this condition is that, in this way, the dynamically-generated hadronic scale $\Lambda_{\scriptscriptstyle{\mathrm{QCD}}}$ is kept fixed, since the QCD $\beta$-function is just a function of $\lambda$ for $N_{\scriptscriptstyle{\mathrm{f}}}/N=0$. The physical mechanism underlying the U(1)A puzzle is identified by noting a peculiarity of the expansion of physical observables in powers of $1/N$ tHooft:1973alw : quark contributions are suppressed by a factor $1/N$ with respect to gluonic ones, so that in this expansion only gluons contribute to the zeroth order (i. e., in the strict large-$N$ limit), while quarks contributions start from the $O(1/N)$ term. However, this is at odds with the fact that $\theta$-dependence should disappear in the presence of massless quark, i. e., in the chiral limit, because of the chiral anomaly: this fact should be valid independently of $N$, however the anomaly is a quark contribution which should be suppressed at large $N$.
In order to solve this problem we can hypothesize that, in the large-$N$ limit, there is a flavor-singlet meson with axial quantum numbers, the $\eta^{\prime}$ meson, whose mass squared vanishes at large-$N$ as $1/N$, so that its propagator at zero momentum is of $O(N)$: that makes $\eta^{\prime}$’s contribution leading order in the $1/N$ expansion, thus able to cancel the gluons’ one . In particular, applying this argument to the topological charge density correlation function, one obtains $m^{2}_{\eta^{\prime}}=2N_{{\scriptscriptstyle{\mathrm{f}}}}\,{\chi^{\infty}_{{}_{\scriptscriptstyle{\mathrm{YM}}}}}/{F_{\pi}^{2}}$ in the chiral limit Witten:1979vv , or the celebrated Witten–Veneziano equation when the finite quark masses are taken into account Veneziano:1979ec ; Witten:1980sp ; DiVecchia:1980yfw :
$\chi_{{}_{\scriptscriptstyle{\mathrm{YM}}}}^{\infty}=\frac{F_{\pi}^{2}}{2N_{\scriptscriptstyle{\mathrm{f}}}}\left(m_{\eta^{\prime}}^{2}+m_{\eta}^{2}-2m_{K}^{2}\right)\,.$ (48)
Here, $\chi_{{}_{\scriptscriptstyle{\mathrm{YM}}}}^{\infty}$ is the topological susceptibility computed in the pure gauge theory for $N\to\infty$: the right-hand side of Eq. (48) is of order zero in the $1/N$ expansion, so the Witten–Veneziano mechanism requires a finite non-zero limit for the pure gauge susceptibility in the $N\to\infty$ limit which, taking into account phenomenological values of the meson masses and of the pion decay constant $F_{\pi}$, turns to be $\chi_{{}_{\scriptscriptstyle{\mathrm{YM}}}}^{\infty}\simeq\left(180~\mathrm{MeV}\right)^{4}$.
A non-zero $\chi_{{}_{\scriptscriptstyle{\mathrm{YM}}}}$ in the $N\to\infty$ limit is at odds with the DIGA model, which predicts no $\theta$-dependence at all, hence $\chi=0$, in the same limit: this is clear from the fact the instanton action $8\pi^{2}/g^{2}=8N\pi^{2}/\lambda$ diverges as $N\to\infty$, so that the 1-instanton weight vanishes, and is also visible from the asymptotic behavior of $\chi$ in Eq. (47). However, we already know that the DIGA model has few hopes to be consistent but in the high temperature regime, hence we do not expect it to match low-energy phenomenological predictions. The point is then whether one can model and predict the low-temperature $\theta$-dependence in the large-$N$ limit by other means: this can indeed be done on quite general grounds Witten:1980sp ; Witten:1998uka .
On one hand, one expects the free energy density to be proportional to the number of degrees of freedom, which is $\sim N^{2}$ in the large-$N$ ’t Hooft limit, since we have $N^{2}-1$ gluons for SU($N$) and fermions become irrelavant. On the other hand, one expects the relevant variable to be $\bar{\theta}=\theta/N$ instead of $\theta$, as can be argued observing that, rescaling the fields by the coupling constant $A^{a}_{\mu}\to\frac{1}{g}A^{a}_{\mu}$ to expose the ’t Hooft coupling, the $\theta$-dependent Yang–Mills action can be expressed at large-$N$ as:
$\mathcal{S}^{{\scriptscriptstyle{(\mathrm{YM})}}}_{{\scriptscriptstyle{\mathrm{E}}}}(\theta)=\frac{N}{\lambda}\left[\frac{1}{4}\int\mathrm{d}^{4}x\,F^{a}_{\mu\nu}(x)F^{a}_{\mu\nu}(x)-\mathrm{i}\lambda\frac{\theta}{N}Q\right]\,.$ (49)
Putting these things together, one obtains the following prediction for the scaling of the free energy density at large-$N$:
$F(T,\theta,N)\underset{N\,\to\,\infty}{\sim}N^{2}\bar{F}\left(T,\frac{\theta}{N}\right)\,.$ (50)
According to the general parameterization in Eq. (41), the scaling function $\bar{F}$ can be written as
$\bar{F}(T,\bar{\theta})\underset{N\,\to\,\infty}{\sim}\frac{1}{2}\chi(T)\bar{\theta}^{2}\left(1+\sum_{n\,=\,1}^{\infty}b_{2n}(T)N^{2n}\bar{\theta}^{2n}\right)\,.$ (51)
It is useful to stress that this large-$N$ prediction is not necessarily at odds with the DIGA model: indeed, they are in perfect agreement if $\chi=0$, meaning no $\theta$-dependence at all. However, if $\chi\sim\mathcal{O}(N^{0})$ as required by the Witten–Veneziano mechanism, the large-$N$ argument leads to a non-trivial prediction, which differs from DIGA substantially. In this case, indeed, by matching the largeN-N limit for the scaling function, hence finite coefficients for the expansion in terms of $\bar{\theta}$, one can infer the following scaling behavior for the $b_{2n}$ coefficients:
$b_{2n}\underset{N\,\to\,\infty}{\sim}\frac{1}{N^{2n}}\,$, (52)
so that, in the large-$N$ limit only the leading $\sim\mathcal{O}(\theta^{2})$ dependence survives. In other QCD-like theories characterized by a non-trivial $\theta$-dependence, like the two-dimensional $\mathrm{CP}^{N-1}$ models, similar large-$N$ arguments lead to more quantitative predictions for both $\chi$ and the $b_{2n}$ coefficients DAdda:1978vbw ; Luscher:1978rn ; Witten:1978bc ; Campostrini:1991kv ; Campostrini:1991tw ; DelDebbio:2006yuf ; Rossi:2016uce ; Bonati:2016tvi ; Sugeno:2025exv , essentially because in vector (instead of matrix) models large-$N$ computations can be explicitly carried out.
However, the semi-quantitative prediction in Eq. (52) has already quite interesting implications. In particular, in order for the analytic dependence on the scaling variable $\theta/N$ to co-exist with the expected $2\pi$-periodicity in $\theta$, the vacuum energy must have non-analytic points where level crossing of different branches occur. Thus, previous arguments suggests the following shape for $F(T,\theta)$ in the large-$N$ limit:
$F(T,\theta)=\frac{1}{2}\chi(T)\,N^{2}\min_{k\,\in\,\mathbb{Z}}\left(\frac{\theta+2\pi k}{N}\right)^{2},$ (53)
which has cusps in $\theta=\pi+2\pi k$. This non-analyticity suggests that the CP symmetry is spontaneously broken at $\theta=\pi$ in the large-$N$ limit, as far as $\chi(T)\neq 0$.
It is interesting to notice that one can try to interpret the different kinds of $\theta$-dependence which are predicted by the DIGA model on one side, and by large-$N$ arguments at low temperature on the other side, in terms of the different degrees of freedom which are thought to be relevant in the two cases. In the DIGA model, the smooth analytic dependence on $\theta$ emerges by assuming that the relevant objects are instantons and anti-instantons, which carry an integer topological charge. In the non-interacting hypothesis, in particular, one obtains $F(T,\theta)\propto(1-\cos\theta)$; one could actually modify such hypothesis by adding interactions, like when moving from a perfect to a real gas in the virial expansion approach, but still obtain a smooth function of $\theta$ if the interaction is weak enough.
It is then tempting to interpret the smooth dependence on $\theta/N$ (instead of $\theta$) which is found in the large-$N$ approach as evidence that, in this case, the relevant objects carry a fractional topological charge, in particular in units of $1/N$. Indeed such objects, sometimes called instanton-quarks, and their possible role in the confined phase have been conjectured since long Fateev:1979dc ; Berg:1979uq ; Belavin:1979fb ; Gonzalez-Arroyo:1995ynx ; Gonzalez-Arroyo:1995isl ; Kraan:1998sn ; Zhitnitsky:2006sr ; Diakonov:2007nv ; Unsal:2008ch ; Parnachev:2008fy ; Zhitnitsky:2008ha ; Gorsky:2009me ; Nair:2022yqi (see also the recent review Gonzalez-Arroyo:2023kqv and references therein). In this perspective, a non-perturbative determination of $\theta$-dependence can provide information about their dynamics: for instance, if the large-$N$ QCD vacuum could be described in terms of a non-interacting ensemble of such objects (like for instantons in the DIGA model), then one would expect $F(T,\theta)\propto[1-\cos(\theta/N)]$, hence $b_{2}=-1/(12N^{2})$. In the following we will compare this possibility with present lattice results about $b_{2}$ in the large-$N$ limit.
Another important aspect of the solution to the U(1)A provided by the Witten–Veneziano mechanism, is that the chiral and the large-$N$ limit do not commute. Indeed:
$\lim_{m\to 0}\lim_{N\to\infty}\,\chi$ $=$ $\chi_{{}_{\scriptscriptstyle{\mathrm{YM}}}}^{\infty}\neq 0\,$, (54)
$\lim_{N\to\infty}\lim_{m\to 0}\,\chi$ $=$ $0\,.$ (55)
So, let us now discuss the predictions that can be obtained in the opposite limit, $m\to 0$ at fixed $N$, which represents important features of QCD at the physical point, given the small masses of the light quarks.
[…]
$\mathcal{L}_{{\scriptscriptstyle{\mathrm{eff}}}}(\theta)[U]=\frac{F_{\pi}^{2}}{4}\mathrm{Tr}\left[\partial_{\mu}U^{{\dagger}}\partial^{\mu}U\right]+\frac{\Sigma}{2}\mathrm{Tr}\left[\mathcal{M}(\theta)U+\mathcal{M}^{\dagger}(\theta)U^{\dagger}\right]\,,\qquad\mathcal{M}(\theta)\equiv\mathrm{e}^{-\mathrm{i}\theta/2}\mathrm{diag}(m_{u},m_{d})\,$, (56)
[…]
At $T=0$, the free energy density receives contribution only from the vacuum energy, which can be estimated by minimizing the potential for the field $U$ appearing in $\mathcal{L}_{{\scriptscriptstyle{\mathrm{eff}}}}(\theta)$, i. e., $V(\theta)=\Sigma\,\Re\left[\mathcal{M}(\theta)U\right]$. That leads to the following expression DiVecchia:1980yfw (see also Refs. GrillidiCortona:2015jxo ; Luciano:2018pbj ):
$F(T=0,\theta)=-F_{\pi}^{2}\left[m_{\pi}^{2}(\theta)-m_{\pi}^{2}\right]\ ;\ \ \ \ \ \ \ m_{\pi}^{2}=\frac{\Sigma}{F_{\pi}^{2}}(1+z)m_{u}\ ;\ \ \ \ \ \ \ z=\frac{m_{u}}{m_{d}}\,.$ (57)
where $m_{\pi}(\theta)$ is the $\theta$-dependent pion mass:
$m_{\pi}^{2}(\theta)=m_{\pi}^{2}\sqrt{1-\frac{4z}{1+z}\sin^{2}\left(\frac{\theta}{2}\right)}\,.$ (58)
Comparing this expression with the general parameterization in Eq. (41) one finds, for the first two terms in the $\theta$ expansion:
$\chi=\frac{z}{1+z}m_{\pi}^{2}F_{\pi}^{2}=\Sigma\left(\frac{1}{m_{u}}+\frac{1}{m_{d}}\right)^{-1}\,,\qquad b_{2}=-\frac{1}{12}\frac{1+z^{3}}{(1+z)^{3}}\,.$ (59)
Therefore, as $m_{u}=m_{d}=m\to 0$, the topological susceptibility vanishes linearly with the quark mass, $\chi\sim m\to 0$, so that the theory becomes $\theta$-independent as expected, while $b_{2}$, as well as the other $b_{2n}$ coefficients, tends to a constant. It is interesting to notice that, for $m_{u}=m_{d}$ ($z=1$), this constant is smaller than the DIGA prediction by a factor 4, so the fact that $\chi\to 0$ cannot be simply interpreted in terms of a more and more dilute gas of integer charged topological objects; however, remarkably, the DIGA prediction is approached when $z\to 0$, i. e. when $m_{u}\to 0$ at fixed $m_{d}$, as can be seen explicitly also be Taylor expanding the dependence of $m_{\pi}(\theta)$ in Eq. (58) and substituting it in Eq. (57), which leads to $F(T,\theta)\propto(1-\cos\theta)$.
[…]
$\chi$ $=\left(75.5(5)~\mathrm{MeV}\right)^{4},\quad b_{2}$ $=-0.29(2),\quad z=0.48\,$, (60)
$\chi$ $=\left(77.8(4)~\mathrm{MeV}\right)^{4},\quad b_{2}$ $=-0.22(1),\quad z=1\,.$ (61)
For the topological susceptibility, also a N2LO calculation is available in the literature; however, the N2LO contribution accidentally cancels with $O(\alpha_{\rm em})$ LO corrections from Quantum Electrodynamics, thus practically not altering these estimates Gorghetto:2018ocs . Finally, one can extend the analysis also to the finite temperature case, at least as far as $T\lesssim T_{\rm pc}$: for this we refer the reader to Ref. GrillidiCortona:2015jxo .
3 The Lattice QCD approach: basics and results
The purpose of this section is to provide a brief introduction to the formulation of Lattice Gauge Theories and to the main issues and results regarding the study of $\theta$-dependence by their means. Of course, any reasonable level of detail about the general formulation is out of scope here, for that we refer the reader to some standard textbooks on the subject Creutz:1983njd ; Rothe:1992nt ; Montvay:1994cy ; Gattringer:2010zz .
The most standard formulation is that on a hypercubic, isotropic lattice, with lattice spacing $a$; gauge fields are represented by the elementary parallel transporters connecting adjacent lattice sites, which are called gauge links and belong to the gauge group, while quarks live on lattice sites. Gluonic gauge invariant observables are written in terms of traces of gauge link products over closed loops: the most elementary is the so-called plaquette, which is related to a possible discretization of the pure gauge action Wilson:1974sk . The covariant derivative in the quark action can also be easily discretized in terms of gauge links, however some difficulties are met to maintain at the same time locality, positivity, the correct number of quark degrees of freedom, and a correct implementation of chiral symmetry Karsten:1980wd ; Nielsen:1981hk . The continuum limit is usually well defined: if $N_{{\scriptscriptstyle{\mathrm{f}}}}$ is not too large, asymptotic freedom guarantees the presence of a critical point for a vanishing value of the bare gauge coupling $g$, where all correlation lengths diverge and the system loses memory of discretization; the dependence of the lattice spacing $a$ on $g$ is known by perturbation theory for $g$ small enough, and can be determined non-perturbatively, i. e., by lattice simulations, otherwise. If dynamical quarks are present, their masses should be tuned with $g$, along the so-called line of constant physics (LCP), in order to maintain physical observables (e. g., the hadron masses) unchanged as the continuum limit is approached.
Once discretized over a finite space-time volume, in order to deal with a finite number of degrees of freedom, the partition function can be sampled by Markov Chain Monte Carlo algorithms (if no sign problem is present), achieving a statistical error on physical observables which decreases as the square root of the statistics. Systematic errors due to discretization can in principle be completely removed by determining observables for several values of the lattice spacings, then performing a continuum extrapolation. The needed computational effort increases as $a$ decreases, both since the number of degrees of freedom increases, and because the auto-correlation times of the sampling algorithms grow as the continuum limit is approached (critical slowing down). The presence of the fermion determinant, which is non-local in the gauge fields, makes numerical simulations of full QCD much more demanding in terms of computational power (even orders of magnitude) with respect to pure gauge simulations, and this is especially true in the presence of light quarks and in the low-temperature phase, because of the accumulation of small eigenvalues of the Dirac operator due to chiral symmetry breaking Banks:1979yr . The adoption of fermion discretizations which optimize their chiral properties (Ginsparg–Wilson fermions Ginsparg:1981bj , in particular domain wall Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky or overlap Neuberger:1997fp ; Neuberger:1998wv ; Luscher:1998pqa ) makes the needed effort much larger.
Topology is a property of continuum fields and of their homotopy classes, therefore, strictly speaking, topology is not well defined on the lattice. However, one expects the concept of topology to be recovered as the continuum limit is approached, so that the lattice investigation of $\theta$-dependence must rely on proper observables, capable of catching the emergence of topology, with a well behaved continuum limit.
One possibility is to interpolate lattice fields by continuum ones, then defining the winding number on the interpolation: that leads to the so-called geometric definitions Berg:1981er ; Luscher:1981zq ; Woit:1983tq ; Phillips:1986qd , which are typically non-polynomial functions of the lattice fields. The lack of well defined homotopy classes then reflects in the presence of lattice configurations with a large but finite action, typically associated to large fluctuations at the lattice spacing scale, for which the interpolation becomes ambiguous: such configurations, despite their large action, can in principle proliferate as the lattice spacing is decreased, for entropic reasons, thus obscuring the correct approach to the continuum limit Berg:1981nw ; Luscher:1981tq . The alternative is to adopt a field-theoretic discretization of $q(x)$, which is polynomial in the gauge fields: however, in this case the discretized topological charge is non-integer and renormalizations appear at various levels, e. g., a multiplicative renormalization Campostrini:1988cy ; Campostrini:1989dh for the topological charge itself, and also additive renormalizations for the susceptibility and the higher order cumulants DiVecchia:1981aev ; DiVecchia:1981hh ; Campostrini:1989dh ; DElia:2003zne , with the latter which can be divergent and thus obscure the continuum limit.
One can also rely on the index theorem and look at the spectrum of a discretized Dirac operator, in particular at its zero modes. Standard lattice discretizations of the Dirac operator do not possess exact zero modes with well defined chirality: in this case one can go on and provide some definition based on the spectrum and the corresponding eigenvectors, which is however subject to renormalizations as well Smit:1987zh ; Smit:1987jh . On the other hand, when dealing with lattice Dirac operators with exact zero-modes (like those satisfying the Ginsparg–Wilson relation), one recovers an integer valued definition of the winding number Giusti:2001xh ; Giusti:2004qd , however ambiguities may emerge also in this case, see e. g. DelDebbio:2003rn ; DelDebbio:2004ns .
A general, successful strategy, is to get rid of field fluctuations at the lattice spacing scale, which are responsible for the renormalizations and ambiguities mentioned above. For gluonic definitions, this is achieved by smoothing techniques, like cooling Berg:1981nw ; Iwasaki:1983bv ; Teper:1985rb ; Ilgenfritz:1985dz ; Campostrini:1989dh , the gradient flow Narayanan:2006rf ; Luscher:2009eq ; Luscher:2010iy ; Luscher:2011bx , or other similar algorithms, all giving comparable results Bonati:2014tqa ; Alexandrou:2017hqw , which drive lattice configurations obtained by Monte Carlo sampling towards minima of the discretized action. For fermionic definitions, one achieves the goal of removing additive renormalizations by projecting observables onto the space generated by the eigenvectors of the lowest part of the Dirac spectrum, thus removing contributions from the ultraviolet (UV) part of it Luscher:2004fu ; Giusti:2008vb ; Luscher:2010ik ; Bonanno:2019xhg . The strategy relies on the fact that the physical scale of topological fluctuations becomes well separated, as the continuum is approached, from the UV scale of the unphysical modes which are smoothed away; moreover, in the continuum topology becomes stable under smoothing.
Topological properties of the gauge ensemble also affect the efficiency of Monte Carlo sampling. A general phenomenon, affecting all theories with a topological classification, is the so-called topological freezing: as the continuum limit is approached, the emergence of infinite action barriers, separating the different topological sectors, makes most Monte Carlo algorithms unable to efficiently move from one sector to the other, thus leading to a severe loss of ergodicity, usually much worse than standard critical slowing down Alles:1996vn ; DelDebbio:2002xa ; DelDebbio:2004xh ; Schaefer:2010hu ; Bonati:2017woi .
Further issues emerge for theories with light fermions. In this case, the fermion determinant should suppress the contribution from configurations with non-zero winding number, because of the associated zero-modes of the Dirac operator, thus leading to the expected suppression of $\theta$-dependence. However, for discretized Dirac operators usually adopted for Monte Carlo sampling, which do not have exact chiral properties, zero-modes are clearly identified only quite close to the continuum limit, so that the observed $\theta$-dependence can be misleading until sufficiently small lattice spacings are reached, typically smaller than those needed for pure gauge theories. On the other hand, the need to reach smaller lattice spacings makes the impact of topological freezing more destructive.
Figure 1: Left: extrapolation of $\chi/\sigma^{2}$, where $\sigma$ is the string tension, towards the large-$N$ limit, using the fit function $\chi/\sigma^{2}=\bar{\chi}/\sigma^{2}+k/N^{2}$; the best fit yields $\bar{\chi}/\sigma^{2}=0.0199(10)$ and $k=0.082(17)$. (source: Ref. Bonanno:2020hht ). Right: Extrapolation of $b_{2}$ towards the large-$N$ limit (source: Ref. Bonanno:2020hht ); a best fit according to $b_{2}=\bar{b}_{2}/N^{c}$ with $c=2$ yields $\bar{b}_{2}=-0.193(10)$ (dashed line), and is stable within errors by making $c$ a free parameter ($c=2.17(26)$, solid line) or by adding $1/N^{4}$ corrections (dotted line).
Let us now go to a brief review of available lattice results. The Witten–Veneziano (WV) formula was the first available prediction about the topological properties of Yang–Mills theories: it was based on large-$N$ considerations and phenomenological inputs, however its explicit confirmation could only be based on some non-perturbative approach, based on first principles. Remarkably, the first pioneer lattice simulations of non-abelian gauge theories Creutz:1979dw ; Creutz:1980zw were performed soon after the proposal of the WV mechanism. Therefore, it is quite natural that the topological susceptibility of SU($N$) pure gauge theories became soon one of the most hunted prey for the first lattice pioneers.
The first attempts DiVecchia:1981aev ; DiVecchia:1981hh were based on a simple gluonic discretization of the topological charge density operator, representing $F_{\mu\nu}$ in terms of an open plaquette, then subtracting an additive renormalization for $\chi$ based on perturbative computations. Results provided a non-zero value for $\chi$ in the SU(2) pure gauge theory, however about 2 orders of magnitude smaller than predicted by the WV formula. Later studies clarified that the problem was in a missing multiplicative renormalization constant for the discretized topological charge density Campostrini:1988cy ; Campostrini:1989dh : after taking that into account, later determinations led to results confirming the WV mechanism also quantitatively, both for SU(2) and SU(3) pure gauge theories Alles:1996nm ; Alles:1997qe . The value of $\chi$ for pure gauge SU(3) theory has been verified by many other studies, adopting gluonic definitions in combination with smoothing techniques Durr:2006ky ; Ce:2015qha ; Bonati:2015sqt ; Athenodorou:2020ani ; Athenodorou:2021qvs ; Bonanno:2023ple ; Durr:2025qtq , or fermionic definitions DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg , achieving an increasing control over the continuum extrapolation and other possible systematic effects. Recently, also the so-called slope of the topological susceptibility, related to the leading order momentum dependence of the two point function of $q(x)$, has been determined Bonanno:2023ple .
The main point of the WV mechanism is however also that $\chi$ is finite in the large-$N$ limit; this has been checked explicitly by various lattice investigations Ce:2016awn ; Bonati:2016tvi ; Bonanno:2020hht ; Athenodorou:2021qvs , which have determined $\chi$ for different values of $N$ and extrapolated it for $N\to\infty$: an example, taken from Ref. Bonanno:2020hht , is reported in the left-hand side of Fig. 1. As discussed above, a finite value of $\chi$ for $N=\infty$ fixes, by large-$N$ arguments, the scaling with $N$ of the higher-order coefficients $b_{2n}$ in the Taylor expansion of $F(T,\theta)$ around $\theta=0$. As explicitly shown in Eq. (42), the determination of such coefficients involves the determination of higher order cumulants of the topological charge distribution at $\theta=0$, which represents a quite difficult numerical challenge: indeed, such cumulants represent the deviations from a purely Gaussian behavior for the probability distribution, however, by the central limit theorem, such deviations are less and less visible as the total four-dimensional volume increases, so that a significant increase in statistics is needed, in order to keep a constant statistical error.
For this reason, early determinations of $b_{2}$ for the SU(3) pure gauge theory were affected by statistical errors of the order of 20% or larger DelDebbio:2002xa ; DElia:2003zne ; Giusti:2007tu . A substantial improvement has been achieved by adding to the action an explicit source term coupled to the topological charge, in particular a purely imaginary $\theta$ term, so as to avoid the sign problem. In this way, the information about the $b_{2n}$ coefficients is transferred to the dependence on the source term of lower order cumulants, with a sizable reduction of statistical errors Panagopoulos:2011rb ; Bonati:2015sqt , bringing the uncertainty for SU(3) well below the 10% level. Further studies have looked at $b_{2}$ also for larger values of $N$ Bonati:2016tvi ; Bonanno:2020hht , obtaining results well compatible with the predicted $1/N^{2}$ scaling: an example, taken from Ref. Bonanno:2020hht , is reported in the right-hand side of Fig. 1. For the same reason, the determination of the other $b_{2n}$ coefficients turns out to be more and more challenging as $n$ increases: as a matter of fact, only $b_{4}$ has been determined, and only for the SU(2) pure gauge theory Bonanno:2018xtd .
As discussed above, for SU($N$) gauge theories large-$N$ arguments can only predict the scaling with $N$, however the measured values of the pre-factor can teach something about the topological configurations dominating the path-integral. For instance, the fact that $b_{2}\simeq-0.19/N^{2}$ (see Fig. 1) indicates that one expects at least sizable corrections to the simple picture of a non-interacting gas of isolated objects with fractional charge $1/N$, which would lead instead to $b_{2}=-0.08333/N^{2}$. This is consistent with the picture emerging from various lattice investigations, which show that, in the low-temperature phase of Yang–Mills theories, the topological charge density is typically delocalized in the dominant gauge configurations (see, e. g., Refs. Buividovich:2011cv ; Ilgenfritz:2007xu ; Alexandru:2005bn ).
[…]
In the framework of a numerical computation of the QCD path-integral, dynamical fermions are encoded in the fermion determinant, which for light masses will act as a suppression term for the statistical weight of gauge configurations with zero-modes of the Dirac operator, i. e. with non-trivial topology. Therefore, a failure of the discretized Dirac operator to correctly identify such zero-modes, will result in a incorrect distribution of the topological charge. While the commonly used lattice discretizations are expected to eventually recover the correct chiral properties close enough to the continuum limit, this effect will typically add to other systematics related to discretization, making the approach to the continuum limit slower. For this reason, as well as for the intrinsic efforts required by simulations with dynamical fermions, reliable investigations of $\theta$-dependence at the physical point, or approaching the chiral limit, have been available only recently; by now, several consistent studies exist Bonati:2015vqz ; Borsanyi:2016ksw ; Alexandrou:2017bzk ; Athenodorou:2022aay ; Bonanno:2023xkg .
[…]
Figure 2: Left: continuum extrapolation of $\chi^{1/4}$ at $T=0$, for $N_{{\scriptscriptstyle{\mathrm{f}}}}=2+1$ QCD at the physical point (discretized via staggered fermions) and various definitions of topological observables; SP stands for Spectral Projectors, with $M/m_{s}$ being the cutoff on the Dirac spectrum used for the projection; figure adapted from Ref. Athenodorou:2022aay . Right: for the same theory and discretization, continuum extrapolated values of $\chi$ for different values of the light-to-strange quark mass ratio, and extrapolation to the chiral limit; figure taken from Ref. Bonanno:2023xkg .
4 High temperature: $\theta$-dependence across the phase transition
As we have discussed above, one expects that the DIGA should give the correct description of $\theta$-dependence at least for asymptotically high temperatures, because of the natural infrared cut-off on large size instantons; however, how high should the temperature be for the DIGA regime to sets in, is in principle not known. On the other hand, we know that in the large-$N$ limit DIGA predicts $\chi=0$; therefore, if $\chi\neq 0$ in the same limit in the zero-$T$ regime, as predicted by the WV mechanism and verified by lattice simulations, there must be a non-analytic point between the two regimes, i. e. a phase transition, at least in the $N\to\infty$ limit. It is then a quite natural expectation that this transition should coincide with the only other finite $T$ phase transition which is known for SU($N$) gauge theories in the large-$N$ limit, namely the deconfinement transition. One can also further conjecture that a change in $\theta$-dependence should persist also for finite values of $N$, still taking place around the deconfinement transition. Remarkably, this is exactly what has been observed by lattice simulations, as we briefly review in the following.
Early indications for a suppression of $\chi$ across the deconfinement transition of the SU(2) pure gauge theory Teper:1985gi ; Teper:1985ek were later confirmed both for SU(3) Alles:1996nm and SU(2) Alles:1997qe . In particular, the topological susceptibility appears to be practically independent of $T$ as long as $T<T_{\rm c}$, where $T_{\rm c}$ is the deconfinement transition, then rapidly dropping to small but non-zero values for $T>T_{\rm c}$. The extension of such results to larger values of $N$ DelDebbio:2004vxo then showed that $\chi$ seems indeed to vanish, as $N\to\infty$, for any $T>T_{\rm c}$, as expected from the argument above. Recently, this has been further corroborated by a calculation of the susceptibility at the critical temperature in the deconfinement phase $\chi_{\rm d}^{+}=\chi(T=T_{\rm c}^{+})$ Borsanyi:2022fub ; Bonanno:2023hhp , showing that this quantity vanishes as $\sim\exp(-N)$ approaching the large-$N$ limit Bonanno:2023hhp . This is a strong indication that the DIGA regime indeed sets in, also for finite $N$, as soon as the system crosses deconfinement: further studies of pure gauge topology at finite $T$ have further refined this picture Berkowitz:2015aua ; Borsanyi:2015cka ; Frison:2016vuc ; Kitano:2015fla ; Borsanyi:2021gqg looking also for the asymptotic behavior of $\chi$ as a function of $T$ predicted at the 1-loop level, see Eq. (47), which, as we have stressed above, could set in at temperatures much larger than those at which the diluteness hypothesis starts to be valid.
Actually, the suppression of $\chi$, even if strongly suggestive, cannot be considered by itself as conclusive evidence for the validity of DIGA, i. e. of the hypothesis that the dominant topological contributions to the functional integral can be described as the superposition of a diluted ensemble of non-interacting objects carrying integer topological charge. Indeed, such hypothesis leads to a precise prediction for $F(T,\theta)\propto(1-\cos\theta)$, see Eq. (45), which implies the values of the $b_{2n}$ coefficients reported in Eq. (46), e. g., $b_{2}=-1/12\simeq-0.833$. There are some counterexamples where a vanishing $\chi$ does not correspond to that, like the chiral limit of QCD with non-vanishing $m_{u}/m_{d}$ (see our discussion after Eq. (59), or two-dimensional $\mathrm{CP}^{N-1}$ models, for which in the $N\to\infty$ limit $\chi$ vanishes as $1/N$ DAdda:1978vbw ; Luscher:1978rn ; Witten:1978bc ; Campostrini:1991kv ; Campostrini:1991tw , however the $b_{2n}$ coefficients vanish as well DelDebbio:2006yuf ; Rossi:2016uce ; Bonati:2016tvi ; Sugeno:2025exv .
For this reason, the lattice investigation of $b_{2}$ as a function of $T$ reported in Ref. Bonati:2013tt represented an important consistency check of the change of $\theta$-dependence happening across the deconfining transition, proving that indeed $b_{2}$ goes to the DIGA prediction soon after $T_{\rm c}$, and increasingly faster as $N$ increases. Such results have been later confirmed by other studies Borsanyi:2015cka ; Vig:2021oyt ; Kovacs:2023vzi ; Yamada:2024vsk and extended to other theories characterized by a deconfining transition, like the 4D gauge theory based on the G2 gauge group Bonati:2015uga , thus proving a general link between the change of $\theta$-dependence and deconfinement.
When considering lattice investigations of $b_{2}$ above $T_{\rm c}$, one should actually be careful about the following possible caveat. Since $\chi$ drops rapidly above $T_{\rm c}$, and since lattice simulations are always constrained to finite volumes, it may easily happen that at high enough temperatures (or too small volumes) $\langle Q^{2}\rangle=\mathcal{V}\,\chi\ll 1$. In this situation, the gauge ensemble is essentially composed by configurations with $Q=0$, with only a small fraction $\epsilon$ with $|Q|=1$ and practically none with $|Q|>1$, so that $Z(T,\theta)\propto(1+\epsilon\cos\theta)$ and $F(T,\theta)\propto(1-\cos\theta)$. However, this is clearly a fake DIGA behaviour, driven by finite volume effects, which is quite different from a fair DIGA behaviour in which $\langle Q^{2}\rangle\gtrsim 1$ and nevertheless one has $F(T,\theta)\propto(1-\cos\theta)$ because of the Poissonian distribution of integer charged topological excitations. The lattice determinations of $b_{2}$ across $T_{\rm c}$ are typically not affected by this possible caveat.
Interestingly, the change of $\theta$-dependence across the phase transition drives a corresponding dependence of the critical temperature $T_{\rm c}$ on $\theta$ Sasaki:2011cj ; Unsal:2012zj ; DElia:2012pvq ; Otake:2022bcq . Indeed, $T_{\rm c}$ is fixed by the exact balance between the free energies of the two phases, confined and deconfined, so that, if the dependence of the two free energies on $\theta$ is different, the balance point is expected to change with $\theta$; in particular, since the confined free energy grows more rapidly (because of the large value of $\chi$), the values of $T_{\rm c}$ is expected to drop with $\theta$, with a leading behaviour proportional to $\theta^{2}/N^{2}$ DElia:2012pvq . This has been explicitly verified by lattice simulations performed at imaginary values of $\theta$, first for $N=3$ DElia:2012pvq ; DElia:2013uaf , then also for $N=2$ Yamada:2024pjy and $N>3$ Bonanno:2023hhp .
The introduction of dynamical fermions does not change the picture significantly. One important point is that, in QCD with physical or close-to-physical quark masses, the deconfinement phase transition is substituted by a rapid crossover at a pseudocritical temperature $T_{\rm pc}$, at which the (approximate) chiral symmetry gets restored, with $T_{\rm pc}\simeq 155$ MeV for physical quark masses. Hence, one can in principle expect that also the change in $\theta$-dependence could be spread over a wider range of temperatures, with respect to the pure gauge case. The drop of the topological susceptibility across the crossover has been typically confirmed by various lattice studies over the years Alles:2000cg ; Bonati:2015vqz ; Borsanyi:2016ksw ; Petreczky:2016vrs ; Bonati:2018blm ; Burger:2017xkz ; Athenodorou:2022aay ; Chen:2022fid ; Kotov:2025ilm and proved to extend also to other transitions in the QCD phase diagram, like in the presence of background magnetic fields Brandt:2024gso or for two-color QCD at finite density Alles:2006ea ; Hands:2011hd ; Iida:2019rah ; Lombardo:2021jpn ; Kawaguchi:2023olk ; Iida:2024irv ; Itou:2025vcy , with the possible exception of the low-$T$ and high density region of the phase diagram Iida:2019rah ; Iida:2024irv ; Itou:2025vcy . The interesting questions are whether the drop of $\chi$ corresponds also in this case to the onset of the DIGA regime, and what is the exact behavior of $\chi(T)$ for temperatures substantially larger than $T_{\rm pc}$; the latter question is particularly important for axion cosmology, since various models permit to fix the axion coupling constant $f_{a}$, hence its mass, based on $\chi(T)$ and on the observed dark matter abundance Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Wantz:2009it ; DiLuzio:2020wdo . Interesting information about axion physics, in particular regarding the axion thermal production Notari:2022ffe , can be obtained also from the Euclidean time correlator of the topological charge density, which permits to determine the so-called sphaleron rate Kotov:2018vyl ; Altenkort:2020axj ; BarrosoMancha:2022mbj ; Bonanno:2023ljc ; Bonanno:2023thi .
The determination of $b_{2}$ has shown a clear approach to the DIGA prediction above $T_{\rm pc}$, even if possibly slower than in the pure gauge case Bonati:2015vqz ; Kotov:2025ilm , and also the drop of $\chi(T)$ is compatible with the 1-loop DIGA prediction in Eq. (47), at least for the power-law exponent, already starting from unexpectedly low values of $T$ Borsanyi:2016ksw ; Petreczky:2016vrs ; Bonati:2018blm ; Athenodorou:2022aay ; Kotov:2025ilm . The determination of $\chi(T)$ turns out to be a particularly challenging problem, since various numerical problems meet at the same time: discretization effects induced by the bad chiral properties of the fermion determinant can lead to misleading results (see, e. g., the early results reported in Ref. Bonati:2015vqz ); the need to reach quite small lattice spacings exacerbates problems related to topological freezing; finally, the very small values reached by $\chi$ at high $T$ mean that, even on reasonable volumes, one has $\langle Q^{2}\rangle=\mathcal{V}\,\chi\ll 1$, which makes the statistical accuracy on $\chi$ quite low, in this case one can rely on multicanonical approaches to partially mitigate this problem Bonati:2017woi ; Bonati:2018blm ; Athenodorou:2022aay .
4 Conclusions
In this chapter we introduced the $\theta$-dependence of four dimensional gauge theories and some of its main features. In particular, we discussed several approximation schemes that can be used to get quantitative or, in some cases, semi-quantitative information on the $\theta$-dependence of QCD and QCD-like theories in various regimes. We then presented a short review of recent results obtained by carrying out numerical simulations of the lattice discretized theory.
While $\theta$-dependence provides fundamental insights in several aspects of the low-energy QCD phenomenology, and especially for what concerns the pseudoscalar meson spectrum, it also opens new problems. The most famous of these problems is the strong CP problem, which is related to the experimental absence of CP violations in the strong sector of the standard model, despite the fact that the $\theta$ term explicitly breaks CP. This fact forces the physical value of the $\theta$ angle to be surprisingly small, $|\theta|\lesssim 10^{-10}$.
The axion solution of the strong CP problem is definitely the most investigated one, mainly because of the phenomenological appeal of the axion as a dark matter candidate. Since several axion properties (in particular its mass and self-coupling) can be directly related to $\theta$-dependence, in the last years there has been a renewed interest in the $\theta$-dependence of QCD at high temperature. It is indeed possible to use the high-temperature behavior of the topological susceptibility $\chi(T)$ (more precisely: $\chi(T)$ in the range of temperatures between $\approx$150MeV and OO(10)GeV) to put an upper bound on the axion decay constant. Despite the enormous progress made in recent years, several important questions regarding $\theta$-dependence remain open or at least need to be explored further.
The most important short-term goal is clearly to reach a quantitative consensus regarding the values of $\chi(T)$ for temperatures up to at least a few GeV. As discussed in Sec. 4 a coherent global picture of $\theta$-dependence in the high temperature regime of QCD in by now well established, however quantitative (and in a few rare cases even qualitative) discrepancies exist between the results obtained by different groups, using different lattice discretizations and analysis techniques. Concerning this point it is important to stress once more that computations of $\chi(T)$ at high temperature, carried out using physical light quark masses, are extremely challenging. To carry out QCD computation with physical light quark masses is already a major computational challenge in standard conditions, but topology make things even worst. Different fermionic lattice discretizations indeed present different lattice artifacts, and results obtained by using different approaches have to be properly extrapolated to the continuum limit to be compared with each other. Topological freezing and the slow continuum convergence of the low-laying Dirac spectrum make however very difficult to perform controlled continuum limit extrapolations, and this is likely the main reason for the discrepancies present in the literature. To improve this state of affairs the development of new sampling algorithms, which will allow us to perform simulations at smaller lattice spacings, is of the paramount importance, and indeed this is a very active research area, in which some traditional and well-established techniques are being applied together with more innovative approaches, see, e. g., Vicari:1992jy ; Brower:2003yx ; Aoki:2007ka ; Luscher:2011kk ; LSD:2014yyp ; Mages:2015scv ; Bietenholz:2015rsa ; Laio:2015era ; Bonati:2017woi ; Hasenbusch:2017unr ; Luscher:2017cjh ; Bonanno:2020hht ; Florio:2019nte ; Funcke:2019zna ; Kanwar:2020xzo ; Nicoli:2020njz ; Albandea:2021lvl ; Cossu:2021bgn ; Borsanyi:2021gqg ; Fritzsch:2021klm ; Abbott:2023thq ; Eichhorn:2023uge ; Howarth:2023bwk ; Albandea:2024fui ; Bonanno:2024udh ; Vadacchino:2024lob ; Abe:2024fpt ; Bonanno:2024zyn .
[…]
When discussing lattice result in Sec. 3 we focused on $\theta$ dependence for small values of $\theta$ and at equilibrium, discussing aspects that are, in practice, related to the second and to the fourth derivative of the free energy density $f(T,\theta)$ at $\theta=0$. In this regime, although some delicate points still need to be further investigated, we have an overall rather clear physical picture. Concerning what happens for larger values of $\theta$, and in particular at $\theta=\pi$, there are instead only exploratory studies Kitano:2017jng ; Kitano:2020mfk ; Hirasawa:2024fjt with more reliable results existing only for systems in reduced dimensionality Alles:2007br ; Alles:2014tta ; Sulejmanpasic:2020lyq ; Sugeno:2025exv ; Matsumoto:2025fjb . Direct simulations at real $\theta$ values are prevented by the sign problem, and for large $\theta$ it is difficult to obtain reliable results from imaginary $\theta$ simulations or Taylor expansion at $\theta=0$, due to the difficulties of performing the analytic continuation or to evaluate a large enough number of coefficients of the expansion. Real time results, relevant to describe out of equilibrium processes, also present huge difficulties, mainly (but not only) related to the resolution of the inverse problem required to extract spectral densities. The only results available so far in this context are those regarding the sphaleron rate Kotov:2018vyl ; Altenkort:2020axj ; BarrosoMancha:2022mbj ; Bonanno:2023ljc ; Bonanno:2023thi .
{ack}
[Acknowledgments] The work of C. Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033. This work is also partially supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006).
see also Axions; Axions and Axion-like particles: collider searches; Chiral anomaly; Chiral perturbation theory; CP violation in the quark sector; Effective field theory; Gauge topology, semiclassics and instantons; Large-$N$ expansion; Lattice field theory;

### v2

@@
 $\mathcal{S}_{{\scriptscriptstyle{\mathrm{QCD}}}}\longrightarrow\mathcal{S}_{{\scriptscriptstyle{\mathrm{QCD}}}}+\theta Q\ .$ (2)
-While this topological $\theta$ term would be completely irrelevant at the classical level (being it the integral of a four-divergence), it has actually physical consequences in the quantum theory.
+While this topological $\theta$-term would be completely irrelevant at the classical level (being it the integral of a four-divergence), it has actually physical consequences in the quantum theory.
 Indeed the path-integral which is expected to non-perturbatively define QCD as a Quantum Field Theory can be decomposed into a sum over all the topological sectors (i.
 e.
-the subsets of the gauge field space with fixed values of the topological charge) and the introduction of a $\theta$ term changes the relative weight of the different sectors, thus introducing a non-trivial dependence on this parameter in all physical observables.
+the subsets of the gauge field space with fixed values of the topological charge) and the introduction of a $\theta$-term changes the relative weight of the different sectors, thus introducing a non-trivial dependence on this parameter in all physical observables.
 One of the most fundamental consequences of $\theta$-dependence for Standard Model phenomenology is the explicit violation of the charge conjugation and parity (CP) symmetry it introduces in the strong sector, being $Q$ CP-odd by virtue of the presence of the Levi-Civita tensor.
@@
 1 Why $\theta$-dependence
-The $\theta$-dependence of four dimensional gauge theories is typically introduced using the semiclassical approximation and instantons, see, e.
+The $\theta$-dependence of four-dimensional gauge theories is typically introduced using the semiclassical approximation and instantons, see, e.
 g., ColemanBook ; ShifmanBook ; Weinberg2Book ; ZinnJustinBook .
@@
 e., a quantum particle moving on a circumference.
-This system, despite its extreme simplicity, displays several key features which have analogues in four-dimensional gauge theories Jackiw:1979ur ; SmilgaBook ; SchulmanBook ; Bietenholz:1997kr ; Gaiotto:2017yup ; Bonati:2017woi .
+This system, despite its extreme simplicity, displays several key features which have analogues in four-dimensional gauge theories Jackiw:1979ur ; Rothe:1977bv ; SmilgaBook ; SchulmanBook ; Bietenholz:1997kr ; Gaiotto:2017yup ; Bonati:2017woi .
 If we denote by $\phi$ the angle on the circumference, the Hamiltonian of a quantum particle of mass $m$ moving on a circumference of radius $R$ is
@@
 Since $\mathrm{e}^{\mathrm{i}\theta}$ is the eigenvalue of the operator $U$ which has been diagonalized simultaneously with $H$ we also have
-$_{\theta^{\prime}}\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle_{\theta}=2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta)}_{n}(t_{f}-t_{i})/\hbar}=2\pi\delta(\theta-\theta^{\prime})\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}\ .$ (7)
+$_{\theta^{\prime}}\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle_{\theta}=\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta)}_{n}(t_{f}-t_{i})/\hbar}=\delta(\theta-\theta^{\prime})\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}\ .$ (7)
 We can now use the identity
@@
 to select the value of $\theta$ we are interested in: using Eq. (4) and Eq. (6) we have
-$\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=$ (9)
-$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f}+2\pi Q)\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\langle\phi_{f}+2\pi Q,t_{f}|\phi_{i},t_{i}\rangle=$
-$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{\phi(t_{i})=\phi_{i}}^{\phi(t_{f})=\phi_{f}+2\pi Q}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)=\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}\left(L(\phi,\dot{\phi})-\hbar\frac{\theta}{2\pi}\dot{\phi}\right)\mathrm{d}t\right)\ ,$
-and we thus obtain the Lagrangian
+$\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=$ (9)
+$\qquad=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f}+2\pi Q)\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\langle\phi_{f}+2\pi Q,t_{f}|\phi_{i},t_{i}\rangle=$
+$\qquad=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{\phi(t_{i})=\phi_{i}}^{\phi(t_{f})=\phi_{f}+2\pi Q}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)\propto\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}\left(L(\phi,\dot{\phi})-\hbar\frac{\theta}{2\pi}\dot{\phi}\right)\mathrm{d}t\right)\ ,$
+where in the last step $\phi_{i}$ and $\phi_{f}$ are defined modulo $2\pi$, the path-integral is performed by summing on all path topologies and we used $\exp(-\mathrm{i}\frac{\theta}{2\pi}\int_{t_{i}}^{t_{f}}\dot{\phi}\mathrm{d}t)=\mathrm{e}^{-\mathrm{i}Q\theta}\mathrm{e}^{-\mathrm{i}\frac{\theta}{2\pi}(\phi_{f}-\phi_{i})}$.
+We thus obtain the Lagrangian
 $L_{\theta}=\frac{1}{2}mR^{2}\dot{\phi}^{2}-V(\phi)-\hbar\frac{\theta}{2\pi}\dot{\phi}\ .$ (10)
@@
 e.
-the effect of the $\theta$ term in the quantum theory, there are (at least) two ways.
+the effect of the $\theta$-term in the quantum theory, there are (at least) two ways.
 The first, more direct, one consists in explicitly computing the energy spectrum of the model in the free case $V=0$: it is indeed trivial to verify that in this case the energy spectrum explicitly depends on $\theta$, and it is given by
 $E_{n}^{(\theta)}=\frac{\hbar^{2}}{2mR^{2}}\left(n+\frac{\theta}{2\pi}\right)^{2}\ ,\quad n\in\mathbb{Z}\ .$ (11)
-The second method uses the path-integral formulation: the $\theta$ term contributes to the propagator from the state $|\phi_{i}\rangle$ at time $t_{i}$ to the state $|\phi_{f}\rangle$ at time $t_{f}$ with
+The second method uses the path-integral formulation: the $\theta$-term contributes to the propagator from the state $|\phi_{i}\rangle$ at time $t_{i}$ to the state $|\phi_{f}\rangle$ at time $t_{f}$ with
 $\exp\left(-\frac{\mathrm{i}\theta}{2\pi}\int_{t_{i}}^{\,t_{f}}\dot{\phi}\mathrm{d}t\right)=\exp\left(-\frac{\mathrm{i}\theta}{2\pi}[\phi_{f}(t_{f})-\phi_{i}(t_{i})]\right)\ .$ (12)
-The effect of the $\theta$ term is thus of topological origin, as can be seen by noting that, for fixed $\phi_{i}(t_{i})$ and $\phi_{f}(t_{f})$, the contributions of two different paths have a relative weight $\exp(-\mathrm{i}\theta\Delta Q)$, where $\Delta Q$ is the difference between the winding numbers along the circumference of the two paths.
+The effect of the $\theta$-term is thus of topological origin, as can be seen by noting that, for fixed $\phi_{i}(t_{i})$ and $\phi_{f}(t_{f})$, the contributions of two different paths have a relative weight $\exp(-\mathrm{i}\theta\Delta Q)$, where $\Delta Q$ is the difference between the winding numbers along the circumference of the two paths.
+Note that the winding number of two open (i.e., with $\phi_{f}\neq\phi_{i}$) paths is not well defined, but their difference is.
 For the particular case of the quantum particle moving on a circumference we can also note that $\theta$ enters the Lagrangian $L_{\theta}$ in the same way as a constant vector potential directed along the $\phi$ direction:
@@
 g., BallentineBook ; KonishiPaffutiBook .
-We are now ready to discuss the case of four dimensional gauge theories, considering for the sake of the simplicity the case of the gauge group SU($N$) and using natural units.
+We are now ready to discuss the case of four-dimensional gauge theories, considering for the sake of the simplicity the case of the gauge group SU($N$) and using natural units.
 The Lagrangian density of the pure glue theory is
@@
 e.
-$\Omega(\mathbf{x})\to 1$ sufficiently fast as $|\mathbf{x}|\to\infty$ (we denote by $\mathbf{x}$ the spatial components of $x$, at fixed time).
+$\Omega(\mathbf{x})\to 1$ sufficiently fast as $|\mathbf{x}|\to\infty$ (we denote by $\mathbf{x}$ the spatial components of $x$, at fixed time); this requirement is needed in order to disentangle the local transformations from the global ones, that are associated to the conserved charges.
 Infinitesimal gauge transformations, in particular, have to satisfy $\chi^{a}(\mathbf{x})\to 0$ as $|\textbf{x}|\to\infty$.
@@
 $\Psi[\mathbf{A}+\delta\mathbf{A}]-\Psi[\mathbf{A}]\simeq\int\frac{\delta\Psi[\mathbf{A}]}{\delta A_{j}^{a}}\delta A_{j}^{a}\mathrm{d}^{3}\mathbf{x}=-\int\left(\chi^{a}\partial_{j}\frac{\delta}{\delta A_{j}^{a}}+gf^{abc}\chi^{a}A_{j}^{b}\frac{\delta}{\delta A_{j}^{c}}\right)\Psi[\mathbf{A}]\mathrm{d}^{3}\mathbf{x}\propto\int\chi^{a}(\mathbf{x})G^{a}\Psi[\mathbf{A}]\mathrm{d}^{3}\mathbf{x}\ .$ (19)
-Since $\chi^{a}(\mathbf{x})$ approaches zero for large x, the previous integral is well defined, and we see that the constraint in Eq.
+Since $\chi^{a}(\mathbf{x})$ approaches zero for large x we could integrate by part in the second step of the previous equation, and the final integral is well defined.
+We thus see that the constraint in Eq.
 (16) implements gauge invariance under infinitesimal local gauge transformations.
@@
 We are now almost ready to adapt the reasoning carried out in Eq.
-(9) to the case of four dimensional gauge theories.
-The last ingredient needed is a relation between the winding number $N(\Omega)$ and four dimensional integrals.
-To this purpose we introduce the (non gauge invariant) Chern-Simons current
-$j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\epsilon^{\,\mu\alpha\beta\gamma}\mathrm{Tr}\left(F_{\alpha\beta}A_{\gamma}+\frac{2\mathrm{i}g}{3}A_{\alpha}A_{\beta}A_{\gamma}\right)\ ,$ (23)
+(9) to the case of four-dimensional gauge theories.
+The last ingredient needed is a relation between the winding number $N(\Omega)$ and four-dimensional integrals.
+To this purpose we introduce the (non gauge invariant) Chern–Simons current
+$j_{\scriptscriptstyle{\rm CS}}^{\mu}=\frac{g^{2}}{16\pi^{2}}\epsilon^{\,\mu\alpha\beta\gamma}\mathrm{Tr}\left(F_{\alpha\beta}A_{\gamma}+\frac{2\mathrm{i}g}{3}A_{\alpha}A_{\beta}A_{\gamma}\right)\ ,$ (23)
 its gauge invariant four-divergence (usually called topological charge density)
-$q\equiv\partial_{\mu}j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\mathrm{Tr}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)=\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (24)
+$q\equiv\partial_{\mu}j_{\scriptscriptstyle{\rm CS}}^{\mu}=\frac{g^{2}}{16\pi^{2}}\mathrm{Tr}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)=\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (24)
 and its integral on a fixed time surface
-$W(\mathbf{A})=\int j_{CS}^{0}(t,\textbf{x})\mathrm{d}^{3}\textbf{x}\ .$ (25)
-Using the four dimensional divergence theorem and assuming the spatial components of the Chern-Simons current to be trivial at spatial infinity (e. g. due to some fixed boundary conditions) we have
+$W(\mathbf{A})=\int j_{\scriptscriptstyle{\rm CS}}^{0}(t,\textbf{x})\mathrm{d}^{3}\textbf{x}\ .$ (25)
+Using the four-dimensional divergence theorem and assuming the spatial components of the Chern–Simons current to be trivial at spatial infinity (e. g. due to some fixed boundary conditions) we have
 $\int q(x)\,\mathrm{d}^{4}x=\int(\partial_{0}j^{0}_{CS}+\partial_{k}j^{k}_{CS})\mathrm{d}^{4}x=\int_{t_{i}}^{t_{f}}\mathrm{d}t\,\partial_{0}W(\mathbf{A})=W(\textbf{A}_{f})-W(\textbf{A}_{i})\ ,$ (26)
@@
 We can now repeat almost verbatim Eq. (9):
-$\left.\langle\mathbf{A}_{f},t_{f}|\textbf{A}_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu(\theta-\theta^{\prime})}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=$ (28)
-$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\langle\mathbf{A}_{f}^{\Omega_{1}^{\nu}},t_{f}|\textbf{A}_{i},t_{i}\rangle=$
-$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}^{\Omega_{1}^{\nu}}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\mathcal{L}\right)\propto\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\left[\mathcal{L}+\theta q\right]\right)\ ,$
+$\left.\langle\mathbf{A}_{f},t_{f}|\textbf{A}_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu(\theta-\theta^{\prime})}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=$ (28)
+$=\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\langle\mathbf{A}_{f}^{\Omega_{1}^{\nu}},t_{f}|\textbf{A}_{i},t_{i}\rangle=$
+$=\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}^{\Omega_{1}^{\nu}}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\mathcal{L}\right)\propto\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\left[\mathcal{L}+\theta q\right]\right)\ ,$
 where in the last step we used the gauge invariance of $\mathcal{L}$ under time dependent gauge transformations and
@@
 2 The strong CP problem
-The Lagrangian density of QCD, taking into account also the matter fields and the $\theta$ term, is thus
+The Lagrangian density of QCD, taking into account also the matter fields and the $\theta$-term, is thus
 $\mathcal{L}=-\frac{1}{4}F^{a}_{\mu\nu}F^{a\,\mu\nu}+\sum_{{\scriptscriptstyle{\mathrm{f}}}}\bar{\psi}_{{\scriptscriptstyle{\mathrm{f}}}}(\mathrm{i}\gamma_{\mu}D^{\mu}-m_{{\scriptscriptstyle{\mathrm{f}}}})\psi_{{\scriptscriptstyle{\mathrm{f}}}}+\theta\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (30)
-where $D_{\mu}=\partial_{\mu}-igA_{\mu}^{a}T^{a}$ is the covariant derivative in the fundamental representation of the gauge group and “f” is a flavor index.
-It is well known that the first two terms of this Lagrangian density (just like in Quantum Electrodynamics) are invariant under both parity (P) and charge conjugation (C), however the $\theta$ term explicitly breaks P and CP for generic values of the $\theta$ angle, and this is essentially due to the presence of the Levi-Civita totally antisymmetric tensor.
-If we introduce the three dimensional chromoelectric and chromomagnetic fields in analogy with classical electromagnetism we have indeed that $q\propto\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, and since the chromomagnetic field is even under parity transformation we see that the $\theta$ term is odd under both P and CP.
+where $D_{\mu}=\partial_{\mu}-\mathrm{i}gA_{\mu}^{a}T^{a}$ is the covariant derivative in the fundamental representation of the gauge group and “f” is a flavor index.
+It is well known that the first two terms of this Lagrangian density (just like in Quantum Electrodynamics) are invariant under both parity (P) and charge conjugation (C), however the $\theta$-term explicitly breaks P and CP for generic values of the $\theta$ angle, and this is essentially due to the presence of the Levi-Civita totally antisymmetric tensor.
+If we introduce the three dimensional chromoelectric and chromomagnetic fields in analogy with classical electromagnetism we have indeed that $q\propto\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, and since the chromomagnetic field is even under parity transformation we see that the $\theta$-term is odd under both P and CP.
 Invariance under parity and CP is clearly recovered for $\theta=0$, but also for $\theta=\pi$: since $\theta$ is defined up to $2\pi$, we have indeed $\pi\sim-\pi$.
@@
 To discuss why this is not possible when considering the whole Standard Model we have to introduce one last ingredient: the chiral anomaly Adler:1969gk ; Bell:1969ts .
-If we consider a four dimensional gauge theory coupled to a single flavor of quark, the chiral rotation
+If we consider a four-dimensional gauge theory coupled to a single flavor of quark, the chiral rotation
 $\psi\to\mathrm{e}^{\mathrm{i}\alpha\gamma^{5}}\psi\ ,\quad\bar{\psi}\to\bar{\psi}\mathrm{e}^{\mathrm{i}\alpha\gamma^{5}}$ (34)
@@
 $m\bar{\psi}\psi\to\bar{\psi}m^{\prime}\frac{1+\gamma^{5}}{2}\psi+\bar{\psi}m^{\prime*}\frac{1-\gamma^{5}}{2}\psi\ ,\quad m^{\prime}=m\mathrm{e}^{2\mathrm{i}\alpha}\ ,\quad\theta\to\theta+2\alpha\ .$ (35)
-This relation implies that we can reabsorb $\theta$ in the mass term at the expense of using the complex mass $m\mathrm{e}^{-i\theta}$.
+This relation implies that we can reabsorb $\theta$ in the mass term at the expense of using the complex mass $m\mathrm{e}^{-\mathrm{i}\theta}$.
 Note that this does not change the invariance properties of the theory, since a complex mass is not invariant under P and CP transformation, but gives us the possibility of studying $\theta$-dependence using chiral perturbation theory if light quarks are present (and this is the rationale behind Eq.
@@
 An immediate consequence of this fact is that there is no $\theta$-dependence at all if $m=0$, since in this case we can shift $\theta$ at will without affecting the mass term of the Lagrangian.
-When both the $\theta$ term and a complex mass are present, we can define the physical $\theta$ angle as the combination $\theta_{phys}=\theta-\arg m$, which is invariant under chiral transformations and corresponds to the value of the $\theta$ angle when the fermion mass is real and positive.
+When both the $\theta$-term and a complex mass are present, we can define the physical $\theta$ angle as the combination $\theta_{\scriptscriptstyle{\rm phys}}=\theta-\arg m$, which is invariant under chiral transformations and corresponds to the value of the $\theta$ angle when the fermion mass is real and positive.
 The previous discussion can be generalized to the case of a generic number of flavors with mass term
 $\bar{\psi}M\frac{1+\gamma^{5}}{2}\psi+\bar{\psi}M^{{\dagger}}\frac{1-\gamma^{5}}{2}\psi\ ,$ (36)
-where $M$ is a $N_{f}\times N_{f}$ matrix and the physical $\theta$ angle is in this case given by
-$\theta_{phys}=\theta-\arg\det M\ .$ (37)
-The previous statement regarding the upper limit of the $\theta$ angle should thus more precisely be written as $|\theta_{phys}|\lesssim 10^{-10}$, and the presence of $\theta_{phys}$ shows why we can not simply fix to zero the $\theta$ angle in the Standard Model: the mass matrix originates from the electroweak sector, and since in the electroweak sector P and CP are not conserved, there is no way to fix $\theta_{phys}=0$ using symmetry arguments.
-The fact that $|\theta_{phys}|$ is so small or, equivalently, the absence of P and CP violations in the strong sector of the Standard Model, constitutes the so called strong CP problem.
+where $M$ is a $N_{{\scriptscriptstyle{\mathrm{f}}}}\times N_{{\scriptscriptstyle{\mathrm{f}}}}$ matrix and the physical $\theta$ angle is in this case given by
+$\theta_{\scriptscriptstyle{\rm phys}}=\theta-\arg\det M\ .$ (37)
+The previous statement regarding the upper limit of the $\theta$ angle should thus more precisely be written as $|\theta_{\scriptscriptstyle{\rm phys}}|\lesssim 10^{-10}$, and the presence of $\theta_{\scriptscriptstyle{\rm phys}}$ shows why we can not simply fix to zero the $\theta$ angle in the Standard Model: the mass matrix originates from the electroweak sector, and since in the electroweak sector P and CP are not conserved, there is no way to fix $\theta_{\scriptscriptstyle{\rm phys}}=0$ using symmetry arguments.
+The fact that $|\theta_{\scriptscriptstyle{\rm phys}}|$ is so small or, equivalently, the absence of P and CP violations in the strong sector of the Standard Model, constitutes the so called strong CP problem.
 As noted before there is no $\theta$-dependence at all if at least one quark is massless ($\arg\det M$ is not defined if $M$ is singular).
@@
 3 General aspects of $\theta$-dependence
-As we argued above, within the general framework of the Standard Model one cannot exclude a priori a $\theta$ term in the QCD Lagrangian.
+As we argued above, within the general framework of the Standard Model one cannot exclude a priori a $\theta$-term in the QCD Lagrangian.
 At the same time, the dependence of various properties of strong interactions on $\theta$ plays a significant role by itself, which is relevant to many phenomenological and theoretical issues, like the interpretation of some particular features of the hadronic spectrum, or the characterization of the various phases which are found when considering Yang–Mills theories or QCD at finite temperature.
-Moreover, one of the proposed solutions to the strong CP problem is the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur , which invokes the presence of a new particle, the axion, which has also become a popular candidate for dark matter: since the axion couples to QCD like a $\theta$ term, the study of $\theta$-dependence in QCD turns out to be fundamental in order to predict some properties of this dark matter candidate.
+Moreover, one of the proposed solutions to the strong CP problem is the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur , which invokes the presence of a new particle, the axion, which has also become a popular candidate for dark matter: since the axion couples to QCD like a $\theta$-term, the study of $\theta$-dependence in QCD turns out to be fundamental in order to predict some properties of this dark matter candidate.
 The purpose of this Section is to go more in detail into these interesting aspects of $\theta$-dependence, and to review the knowledge that we have achieved so far.
@@
 It is clear from Eq.
-(38) that, in the Euclidean formulation, the $\theta$ term gives a purely imaginary contribution to the action, making it complex.
+(38) that, in the Euclidean formulation, the $\theta$-term gives a purely imaginary contribution to the action, making it complex.
 This is usually known as then sign problem, since it ruins the positivity of the Euclidean path-integral measure, thus undermining the Monte Carlo approach at its very basis.
 A similar problem appears in other cases, for instance, the fermion determinants themselves becomes complex after adding chemical potentials coupled to quark numbers.
-In this case, its origin can be traced back to the fact that the $\theta$ term involves an odd number of time derivatives (contrary to the standard pure gauge term), hence the appearance of the extra factor $\mathrm{i}$”.
+In this case, its origin can be traced back to the fact that the $\theta$-term involves an odd number of time derivatives (contrary to the standard pure gauge term), hence the appearance of the extra factor $\mathrm{i}$”.
 That also underlies one of the exact results about $\theta$-dependence Vafa:1983tf ; Vafa:1984xg , the so-called Vafa–Witten theorem.
@@
 $F(T,\theta)=-\frac{1}{\mathcal{V}}\log\left[\frac{Z(T,\theta)}{Z(T,0)}\right]$ (40)
-where $\mathcal{V}$ is the 4-dimensional volume.
-The extra factor $\mathrm{i}$” of the Euclidean $\theta$ term implies that the ratio of partion functions can be viewed as the expectation value of a pure phase factor, $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$, which is then bounded by 1 from above, meaning that $F(T,\theta)\geq 0$ with a minimum in $\theta=0$ [^1] The same property holds for the quantum particle moving on a circumference discussed above: in that case, since a non-zero $\theta$ can be interpreted as the presence of a magnetic flux, the property implies that the system is diamagnetic.
+where $\mathcal{V}$ is the four-dimensional volume.
+The extra factor $\mathrm{i}$” of the Euclidean $\theta$-term implies that the ratio of partion functions can be viewed as the expectation value of a pure phase factor, $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$, which is then bounded by 1 from above, meaning that $F(T,\theta)\geq 0$ with a minimum in $\theta=0$ [^1] The same property holds for the quantum particle moving on a circumference discussed above: in that case, since a non-zero $\theta$ can be interpreted as the presence of a magnetic flux, the property implies that the system is diamagnetic.
 That rules out the possibility of spontaneous P breaking at $\theta=0$ Vafa:1984xg , hence $F(T,\theta)$ is an even function of $\theta$ with period $2\pi$ (because $Q\in\mathbb{Z}$) and analytic around $\theta=0$.
@@
 Eq. (40) and the relation $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$ permit to rewrite the expansion coefficients in terms of expectation values computed at $\theta=0$
-$\chi=\frac{1}{\mathcal{V}}\braket{Q^{2}}\ ;\ \ \ \ \ b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\frac{\braket{Q^{2n}}_{c}}{\braket{Q^{2}}}$ (42)
+$\chi=\frac{1}{\mathcal{V}}\braket{Q^{2}}\ ;\ \ \ \ \ b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\frac{\braket{Q^{2n}}_{c}}{\braket{Q^{2}}}\ ,$ (42)
 where $\braket{\cdot}_{c}$ stands for a connected expectation value, e. g., $\braket{Q^{4}}_{c}=\braket{Q^{4}}-3\braket{Q^{2}}^{2}$.
@@
 An attractive solution is provided by the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma , which invokes an extension of the Standard Model with an extra axial U(1) global symmetry, the U(1)PQ Peccei–Quinn symmetry.
-Such symmetry is supposed to be spontaneously broken, implying the possible emergence of a massless Nambu-Goldstone boson, the axion, but also anomalously broken at the same time, leading to a coupling of the axion field $a(x)$ to the topological charge density operator $q(x)$.
+Such symmetry is supposed to be spontaneously broken, implying the possible emergence of a massless Nambu–Goldstone boson, the axion, but also anomalously broken at the same time, leading to a coupling of the axion field $a(x)$ to the topological charge density operator $q(x)$.
 One can built various theories implementing such features, see e.
-g., Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq for early proposals and DiLuzio:2020wdo for a recent review, however, at the level of a low energy effective theory, they all correspond to the addition to the QCD Lagrangian of a massless pseudoscalar with derivative couplings plus a modification of the $\theta$ term as follows:
+g., Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq for early proposals and DiLuzio:2020wdo for a recent review, however, at the level of a low energy effective theory, they all correspond to the addition to the QCD Lagrangian of a massless pseudoscalar with derivative couplings plus a modification of the $\theta$-term as follows:
 $\theta\,q(x)\ \ \to\ \ \left(\frac{a}{f_{a}}+\theta\right)q(x)$ (43)
@@
 However, if $\chi\sim\mathcal{O}(N^{0})$ as required by the Witten–Veneziano mechanism, the large-$N$ argument leads to a non-trivial prediction, which differs from DIGA substantially.
-In this case, indeed, by matching the largeN-N limit for the scaling function, hence finite coefficients for the expansion in terms of $\bar{\theta}$, one can infer the following scaling behavior for the $b_{2n}$ coefficients:
+In this case, indeed, by matching the largeN-N limit for the scaling function, hence finite coefficients for the expansion in terms of $\bar{\theta}$, one can infer the following scaling behavior for the $b_{2n}$ coefficients (for a large-$N$ analysis of QCD $\theta$-dependence using effective chiral Lagrangians see Vonk:2019kwv ; GomezNicola:2019myi ):
 $b_{2n}\underset{N\,\to\,\infty}{\sim}\frac{1}{N^{2n}}\,$, (52)
@@
 The value of $\chi$ for pure gauge SU(3) theory has been verified by many other studies, adopting gluonic definitions in combination with smoothing techniques Durr:2006ky ; Ce:2015qha ; Bonati:2015sqt ; Athenodorou:2020ani ; Athenodorou:2021qvs ; Bonanno:2023ple ; Durr:2025qtq , or fermionic definitions DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg , achieving an increasing control over the continuum extrapolation and other possible systematic effects.
-Recently, also the so-called slope of the topological susceptibility, related to the leading order momentum dependence of the two point function of $q(x)$, has been determined Bonanno:2023ple .
+Recently, also the so-called slope of the topological susceptibility, related to the leading-order momentum-dependence of the two-point function of $q(x)$, has been determined Bonanno:2023ple .
 The main point of the WV mechanism is however also that $\chi$ is finite in the large-$N$ limit; this has been checked explicitly by various lattice investigations Ce:2016awn ; Bonati:2016tvi ; Bonanno:2020hht ; Athenodorou:2021qvs , which have determined $\chi$ for different values of $N$ and extrapolated it for $N\to\infty$: an example, taken from Ref.
@@
 For this reason, early determinations of $b_{2}$ for the SU(3) pure gauge theory were affected by statistical errors of the order of 20% or larger DelDebbio:2002xa ; DElia:2003zne ; Giusti:2007tu .
-A substantial improvement has been achieved by adding to the action an explicit source term coupled to the topological charge, in particular a purely imaginary $\theta$ term, so as to avoid the sign problem.
+A substantial improvement has been achieved by adding to the action an explicit source term coupled to the topological charge, in particular a purely imaginary $\theta$-term, so as to avoid the sign problem.
 In this way, the information about the $b_{2n}$ coefficients is transferred to the dependence on the source term of lower order cumulants, with a sizable reduction of statistical errors Panagopoulos:2011rb ; Bonati:2015sqt , bringing the uncertainty for SU(3) well below the 10% level.
@@
 The extension of such results to larger values of $N$ DelDebbio:2004vxo then showed that $\chi$ seems indeed to vanish, as $N\to\infty$, for any $T>T_{\rm c}$, as expected from the argument above.
-Recently, this has been further corroborated by a calculation of the susceptibility at the critical temperature in the deconfinement phase $\chi_{\rm d}^{+}=\chi(T=T_{\rm c}^{+})$ Borsanyi:2022fub ; Bonanno:2023hhp , showing that this quantity vanishes as $\sim\exp(-N)$ approaching the large-$N$ limit Bonanno:2023hhp .
+Recently, this has been further corroborated by a calculation of the susceptibility at the critical temperature in the deconfined phase $\chi_{\rm d}^{+}=\chi(T=T_{\rm c}^{+})$ Borsanyi:2022fub ; Bonanno:2023hhp , showing that this quantity vanishes as $\sim\exp(-N)$ approaching the large-$N$ limit Bonanno:2023hhp .
 This is a strong indication that the DIGA regime indeed sets in, also for finite $N$, as soon as the system crosses deconfinement: further studies of pure gauge topology at finite $T$ have further refined this picture Berkowitz:2015aua ; Borsanyi:2015cka ; Frison:2016vuc ; Kitano:2015fla ; Borsanyi:2021gqg looking also for the asymptotic behavior of $\chi$ as a function of $T$ predicted at the 1-loop level, see Eq.
@@
 In this situation, the gauge ensemble is essentially composed by configurations with $Q=0$, with only a small fraction $\epsilon$ with $|Q|=1$ and practically none with $|Q|>1$, so that $Z(T,\theta)\propto(1+\epsilon\cos\theta)$ and $F(T,\theta)\propto(1-\cos\theta)$.
-However, this is clearly a fake DIGA behaviour, driven by finite volume effects, which is quite different from a fair DIGA behaviour in which $\langle Q^{2}\rangle\gtrsim 1$ and nevertheless one has $F(T,\theta)\propto(1-\cos\theta)$ because of the Poissonian distribution of integer charged topological excitations.
+However, this is clearly a fake DIGA behaviour, driven by finite volume effects, which is quite different from a fair DIGA behaviour in which $\langle Q^{2}\rangle\gtrsim 1$ and nevertheless one has $F(T,\theta)\propto(1-\cos\theta)$ because of the Poissonian distribution of integer-charged topological excitations.
 The lattice determinations of $b_{2}$ across $T_{\rm c}$ are typically not affected by this possible caveat.
@@
 4 Conclusions
-In this chapter we introduced the $\theta$-dependence of four dimensional gauge theories and some of its main features.
+In this chapter we introduced the $\theta$-dependence of four-dimensional gauge theories and some of its main features.
 In particular, we discussed several approximation schemes that can be used to get quantitative or, in some cases, semi-quantitative information on the $\theta$-dependence of QCD and QCD-like theories in various regimes.
@@
 While $\theta$-dependence provides fundamental insights in several aspects of the low-energy QCD phenomenology, and especially for what concerns the pseudoscalar meson spectrum, it also opens new problems.
-The most famous of these problems is the strong CP problem, which is related to the experimental absence of CP violations in the strong sector of the standard model, despite the fact that the $\theta$ term explicitly breaks CP.
+The most famous of these problems is the strong CP problem, which is related to the experimental absence of CP violations in the strong sector of the Standard Model, despite the fact that the $\theta$-term explicitly breaks CP.
 This fact forces the physical value of the $\theta$ angle to be surprisingly small, $|\theta|\lesssim 10^{-10}$.
@@
 Since several axion properties (in particular its mass and self-coupling) can be directly related to $\theta$-dependence, in the last years there has been a renewed interest in the $\theta$-dependence of QCD at high temperature.
-It is indeed possible to use the high-temperature behavior of the topological susceptibility $\chi(T)$ (more precisely: $\chi(T)$ in the range of temperatures between $\approx$150MeV and OO(10)GeV) to put an upper bound on the axion decay constant.
+It is indeed possible to use the high-temperature behavior of the topological susceptibility $\chi(T)$ (more precisely: $\chi(T)$ in the range of temperatures between $\approx 150$ MeV and $O(10)$ GeV) to put an upper bound on the axion decay constant.
 Despite the enormous progress made in recent years, several important questions regarding $\theta$-dependence remain open or at least need to be explored further.
@@
 When discussing lattice result in Sec.
-3 we focused on $\theta$ dependence for small values of $\theta$ and at equilibrium, discussing aspects that are, in practice, related to the second and to the fourth derivative of the free energy density $f(T,\theta)$ at $\theta=0$.
+3 we focused on $\theta$-dependence for small values of $\theta$ and at equilibrium, discussing aspects that are, in practice, related to the second and to the fourth derivative of the free energy density $f(T,\theta)$ at $\theta=0$.
 In this regime, although some delicate points still need to be further investigated, we have an overall rather clear physical picture.
 Concerning what happens for larger values of $\theta$, and in particular at $\theta=\pi$, there are instead only exploratory studies Kitano:2017jng ; Kitano:2020mfk ; Hirasawa:2024fjt with more reliable results existing only for systems in reduced dimensionality Alles:2007br ; Alles:2014tta ; Sulejmanpasic:2020lyq ; Sugeno:2025exv ; Matsumoto:2025fjb .
-Direct simulations at real $\theta$ values are prevented by the sign problem, and for large $\theta$ it is difficult to obtain reliable results from imaginary $\theta$ simulations or Taylor expansion at $\theta=0$, due to the difficulties of performing the analytic continuation or to evaluate a large enough number of coefficients of the expansion.
+Direct simulations at real $\theta$ values are prevented by the sign problem, and for large $\theta$ it is difficult to obtain reliable results from imaginary-$\theta$ simulations or Taylor expansion at $\theta=0$, due to the difficulties of performing the analytic continuation or to evaluate a large enough number of coefficients of the expansion.
 Real time results, relevant to describe out of equilibrium processes, also present huge difficulties, mainly (but not only) related to the resolution of the inverse problem required to extract spectral densities.
@@
 {ack}
-[Acknowledgments] The work of C.
+[Acknowledgments] We would like to thank B.
+Garbrecht and F.-K.
+Guo for comments.
+The work of C.
 Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033.